### v1

# Regularization by regular noise: a numerical result

Abstract.
We study a singular stochastic equation driven by a regular noise of fractional Brownian type with Hurst index $H\in(1,\infty)\setminus\mathbb{Z}$ and drift coefficient $b\in\mathcal{C}^{\alpha}$, where $\alpha>1-\tfrac{1}{2H}$. The strong well-posedness of this equation was first established in [Ger23], a phenomenon referred to as regularization by regular noise. In this note, we provide a corresponding numerical analysis. Specifically, we show that the Euler–Maruyama approximation $X^{n}$ converges strongly to the unique solution $X$ with rate $n^{-1}$. Furthermore, under the additional assumption $b\in\mathcal{C}^{1}$, we show that $n(X-X^{n})$ converges to a non-trivial limit as $n\to\infty$, thereby confirming that the rate $n^{-1}$ is in fact optimal upper bound for this scheme.
Mathematics Subject Classification (2020): Primary 60H35, 65C30, 60H10; Secondary 60H50, 60L90.
Keywords: Singular SDEs; Strong approximation; Euler scheme; Regularization by noise; Stochastic sewing, Fractional Brownian motion.
1. Introduction
It is known that when $H\in(0,1)$, a $d$-dimensional fractional Brwonian motion (fBM) $B^{H}$ can be defined via the Mandelbrot - van Ness representation [MVN68]:
$B_{t}^{H}:=\int_{-\infty}^{0}(|t-s|^{H-\frac{1}{2}}-|s|^{H-\frac{1}{2}})\mathrm{d}W_{s}+\int_{0}^{t}|t-s|^{H-\frac{1}{2}}\mathrm{d}W_{s},\quad t\geqslant 0,$ (1.1)
[…]
$B_{t}^{H}:=\int_{0\leqslant s_{1}\leqslant\ldots\leqslant s_{\lfloor H\rfloor}\leqslant t}B^{H-\lfloor H\rfloor}_{s_{1}}\mathrm{d}{s_{1}}\ldots\mathrm{d}s_{\lfloor H\rfloor}.$ (1.2)
Due to the possible multiple times of integral in (1.2) we can see clearly that the path of $B^{H}$ for $H>1$ is regular, at least $C^{1}$. This fact yields one of the appealing points of [Ger23] which shows the strong well-posedness of following equation with singular $b$:
$\mathrm{d}X_{t}=b(X_{t})\mathrm{d}t+\mathrm{d}B_{t}^{H},\quad X_{0}=x_{0}\in{\mathbb{R}}^{d},$ (1.3)
[…]
$\left\{\begin{array}[]{cc}\mathrm{d}X_{t}&=\big(b(X_{t})+V_{t}^{\lfloor H\rfloor}\big)\mathrm{d}t,\\ \mathrm{d}V_{t}^{\lfloor H\rfloor}&=V_{t}^{\lfloor H\rfloor-1}\mathrm{d}t,\\ \qquad\qquad\qquad\qquad\ldots,&\\ \mathrm{d}V_{t}^{1}&=\mathrm{d}B_{t}^{H-\lfloor H\rfloor},\end{array}\right$. (1.8)
which shares the similar framework from [CdRHM22, HZZZ24, HLL24, Lin24] but with non-Markvin noise when $H\neq k+\frac{1}{2}$ for any $k\in{\mathbb{N}}$.
In this note our motivation is to provide a numerical result of it by considering its Euler-Maruyama (EM) scheme
$\mathrm{d}X_{t}^{n}=b(X_{k_{n}(t)}^{n})\mathrm{d}t+\mathrm{d}B_{t}^{H},\qquad X_{0}^{n}=x_{0}^{n}\in{\mathbb{R}}^{d}$ (1.9)
with $k_{n}(t)\coloneqq\frac{\lfloor nt\rfloor}{n}$.
When $H\in(0,1)$, [BDG21] has shown the strong convergence rate $n^{-(\frac{1}{2}+\alpha H)\wedge 1+\epsilon}$ of the scheme (1.9) to (1.3), inside which the Girsanov Theroem and Stochastic Sewing Lemma (SSL) [Lê20] play one of the most important roles. Precisely the methodology inside [BDG21] can be roughly summarized as (take $x_{0}=x_{0}^{n}$)
$\|\sup_{t\in[0,1]}|X_{t}-X^{n}_{t}|\|_{L^{p}_{\omega}}\overset{\text{Girsanov}}{\lesssim}\big\|\int_{0}^{1}b(B_{s}^{H})-b(B_{k_{n}(s)}^{H})\mathrm{d}r\big\|_{L^{p}_{\omega}}\overset{\text{SSL}}{\lesssim}n^{-(\frac{1}{2}+\alpha H)\wedge 1+\epsilon}.$
When $H>1$, on one hand, as indicated in [Ger23, GG25] the Girsanov Theorem becomes less helpful; meanwhile PDE tools clearly would not work. Therefore the challenge for showing the convergence of (1.9) to (1.3) lies in two folds comparing with known results: absence of Girsanov and non-Markovianity. Alternatively, although [Ger23] studies the well-posedness only, actually it already hints the possible way to get rid of Girsanov; concerning numerical approximation [BDG24] provides the similar evidence, however it considers the singular SDE driven by an $\alpha$ stable process which is a Markov process. Nevertheless we are able to show convergence rate $n^{-1}$ which is comparable with [BDG21] since $H>1$ here implies $\alpha>1-\frac{1}{2H}>\frac{1}{2}$.
For such obtained convergence rate, naturally we wonder how far it is away from being optimal. Such question on the optimality of EM scheme has been addressed and studied for instance in [MGY23, EMGY25, DGL23a] concerning Brownian noise, [PSS24] for Lévy processes with jumps and [Neu06, NN07, LLT25, HLN16] for fBM with $H<1$. In particular [Neu06] confirms for $H\in(\frac{1}{2},1)$, order $n^{H}$ reaches its optimality of EM scheme (1.9) for $b\in{\mathcal{C}}^{2}$ (twice differentiable) with possible linear growth. Here we complete this result with rate $n^{-1}$ for $H>1$ and less regular $b$ (actually only ${\mathcal{C}}^{1}$). The idea of the proof is straightforward: we show that for $b\in{\mathcal{C}}^{1}$, the following approximation reaches a possibly non-zero limit:
$n(X_{t}-X^{n}_{t})\overset{n\rightarrow\infty}{\rightarrow}c(t)\neq 0,\quad a.s.$
which indicates that for (1.9) the best convergence rate we can expect should be not faster than order $n$. Evidently it also verifies that the rate we obtained actually is optimal.
Lastly, let us mention a few related works that share a similar interest and spirit in the study of numerical approximations. For equations of the type (1.3) with singular drift, convergence results have been established in [NS21, DG20, JM21, BW25] for additive Brownian motion, in [DGL23b, BDG21, GLL25, LL25] for multiplicative Brownian noise, in [BDG24, BWWZ25] for Lévy processes with jumps, and in [BDG21, GHR25] for fractional Brownian motion. A slightly different notion of singularity-namely, piecewise Lipschitz coefficients-has also been investigated, with convergence results obtained in works such as [LS16, MGY20, MGRY25, PSS24]. We must acknowledge that this is only a brief selection from a vast body of literature, and we encourage readers to explore the cited works for further details and insights.
Organization of the paper
In Section˜2 we introduce the necessary notations and main results. Section˜3 collects all of the crucial tools and properties of fBM. We present the central analysis and proof of the main result in Section˜4. In the end we discuss optimality in Section˜5.
2. Preliminaries and main results
2.1. Preliminaries
On finite dimensional vector spaces we always use the Euclidean norm.
For $k\in\mathbb{N}$, $f:\mathbb{R}^{d}\mapsto\mathbb{R}$, denote $\partial_{k}f(x)\coloneqq\frac{\partial f(x)}{\partial x_{k}}$ for $x\in\mathbb{R}^{d}$ and $\nabla f(x)\coloneqq(\partial_{i}f(x))_{1\leqslant i\leqslant d}$, the derivative is understood in the weak sense. For vector-valued $f$ we use the same notation, and $\nabla^{k}f$ is defined via $\nabla(\nabla^{k-1}f)$ iteratively. For a multi-index $k=(k_{1},\ldots,k_{d})\in{\mathbb{N}}^{d}$, denote $\partial^{k}f(x)\coloneqq\frac{\partial^{|k|}f(x)}{\partial x_{k_{1}}\cdots\partial x_{k_{d}}}$. If $k=(0,\ldots,0)$, we use convention $\partial^{k}f=f$. We denote by ${C}^{\infty}_{0}$ (${C}_{p}^{\infty}$, resp.) the set of all continuously infinitely differentiable functions that, along with all of their partial derivatives, are compactly supported (of polynomial growth, resp.).
For $\alpha\in(0,1]$, we set $\mathcal{C}^{\alpha}({\mathbb{R}}^{d})$ to be the space of continuous functions such that
$\|f\|_{\mathcal{C}^{\alpha}}\coloneqq[f]_{\mathcal{C}^{\alpha}}+\sup_{x\in{\mathbb{R}}^{d}}|f(x)|\coloneqq\sup_{x,y\in{\mathbb{R}}^{d},x\neq y}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}+\sup_{x\in{\mathbb{R}}^{d}}|f(x)|<\infty.$
Here, and often below, we write $\mathcal{C}^{\alpha}$ instead of $\mathcal{C}^{\alpha}({\mathbb{R}}^{d})$ for simplicity. For $\alpha\in(0,\infty)$, we define $\mathcal{C}^{\alpha}({\mathbb{R}}^{d})$ the space of all functions $f$ defined on ${\mathbb{R}}^{d}$ having bounded derivatives $\partial^{k}f$ for multi-indices $k\in{\mathbb{N}}^{d}$ with $|k|\leqslant\alpha$ so that
$\|f\|_{\mathcal{C}^{\alpha}}$ $\coloneqq\|f\|_{\mathcal{C}^{\lfloor\alpha\rfloor}}+[f]_{\mathcal{C}^{\alpha}}\coloneqq\sum_{|k|\leqslant\lfloor\alpha\rfloor}\sup_{x\in{\mathbb{R}}^{d}}|\partial^{k}f(x)|+\sum_{k=\lfloor\alpha\rfloor}[\partial^{k}f]_{\mathcal{C}^{\{\alpha\}}}<\infty,$
where $\{\alpha\}:=\alpha-\lfloor\alpha\rfloor$. Note that the $\mathcal{C}^{\alpha}$-norm always includes the supremum of the function. We also denote the space of bounded measurable functions $\mathcal{C}^{0}({\mathbb{R}}^{d})$ with the supremum norm. To be noticed that the functions in $\mathcal{C}^{0}$ do not need to be continuous.
In the following we denote the conditional expectation w.r.t. the $\sigma$-algebras of the filtration $(\mathcal{F}_{t})_{t\geqslant 0}$ as ${\mathbb{E}}^{t}(\cdot)\coloneqq{\mathbb{E}}(\cdot|\mathcal{F}_{t}),t\geqslant 0$, $\|X\|_{L^{p}_{\omega}}:=({\mathbb{E}}|X|^{p})^{\frac{1}{p}}$, $\|X\|_{L^{p}_{\omega}|{\mathcal{F}}_{s}}:=({\mathbb{E}}[|X|^{p}|{\mathcal{F}}_{s}])^{\frac{1}{p}}$.
For $p\in[1,\infty]$, $X\in L^{p}(\Omega,{\mathbb{R}}^{d})$ and $\mathcal{F}_{s}$-measurable ${\mathbb{R}}^{d}$ valued random vector $Y$, we have the following inequalities
$\|{\mathbb{E}}^{s}X\|_{L^{p}_{\omega}}\leqslant\|X\|_{L^{p}_{\omega}}$ (2.1)
and
$\|X-{\mathbb{E}}^{s}X\|_{L^{p}_{\omega}|\mathcal{F}_{s}}\leqslant 2\|X-Y\|_{L^{p}_{\omega}|\mathcal{F}_{s}}\quad a.s.$ (2.2)
Let $f:[0,1]\times\Omega\rightarrow\mathbb{R}^{d}$ be a measurable function adapted to the filtration $(\mathcal{F}_{t})_{t\geqslant 0}$, $\gamma\in(0,1]$, $p\geqslant 2$ and $S,T]\subset[0,1]$. We give the following definitions:
$\|f\|_{C_{p}^{0}[S,T]}\coloneqq\sup_{r\in[S,T]}\|f(r)\|_{L^{p}_{\omega}};$
$[f]_{C_{p}^{\gamma}[S,T]}\coloneqq\sup_{r_{1},r_{2}\in[S,T],r_{1}\neq r_{2}}\frac{\|\partial^{\lfloor\gamma\rfloor}f(r_{1})-\partial^{\lfloor\gamma\rfloor}f(r_{2})\|_{L^{p}_{\omega}}}{|r_{1}-r_{2}|^{\{\gamma\}}};$
$\|f\|_{C_{p}^{\gamma}[S,T]}\coloneqq\|f\|_{C_{p}^{0}[S,T]}+[f]_{C_{p}^{\gamma}[S,T]}.$
If $f$ is an adapted process, we choose $Y$ in (2.2) as the value at $t$ of the Taylor expansion of $f$ at $s$ up to order $\lfloor\gamma\rfloor$ and we obtain
$\|f_{t}-{\mathbb{E}}^{s}f_{t}\|_{L^{p}_{\omega}}\leqslant 2|t-s|^{\gamma}[f]_{{C}^{\gamma}_{p}[s,t]}.$ (2.3)
In proofs, the notation $a\lesssim b$ (respectively $a\asymp b$) abbreviates the existence of $C>0$ such that $a\leqslant Cb$ (respectively $C^{-1}b\leqslant a\leqslant Cb$), such that moreover $C$ depends only on the parameters claimed in the corresponding statement.
2.2. Main results
Our main assumption and result can be stated as follows.
Assumption 2.1.
[…]
Notice that following [Ger23], under the assumption above, there exists a unique strong solution to (1.3). Here is our numerical approximation result for it.
Theorem 2.2.
Let $(X_{t})_{t\in[0,1]},(X_{t}^{n})_{t\in[0,1]}$ be the solutions to (1.3) and (1.9) accordingly. Suppose Assumption˜2.1 holds. Then for every $p>1$, we have
$\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[0,1]}\leqslant N|x_{0}-x_{0}^{n}|+Nn^{-1},$ (2.4)
where $N=N(p,d,\alpha,H,\|b\|_{{\mathcal{C}}^{\alpha}})$.
Remark 2.3.
As an easy application of Kolmogorov continuity criteria we can also conclude from (2.4) the following:
$\big\|\sup_{t\in[0,1]}|X_{t}-X_{t}^{n}|\big\|_{L_{\omega}^{p}}\leqslant N|x_{0}-x_{0}^{n}|+Nn^{-1}.$
Idea of the analysis. Its detailed proof will be given in Section˜4. Here we only outline the rough idea.
Observe that for any $p\geqslant 1$, for $0\leqslant s\leqslant t\leqslant 1$
$\big\|(X_{t}-X_{t}^{n})-(X_{s}-X_{s}^{n})|\big\|_{L^{p}_{\omega}}=\big\|(\varphi_{t}-\varphi_{t}^{n})-(\varphi_{s}-\varphi_{s}^{n})\big\|_{L^{p}_{\omega}},$
where
$\varphi_{t}:=X_{t}-B_{t}^{H}=\int_{0}^{t}b(X_{r})\mathrm{d}r=\int_{0}^{t}b(\varphi_{r}+B_{r}^{H})\mathrm{d}r,$
$\varphi_{t}^{n}:=X_{t}^{n}-B_{t}^{H}=\int_{0}^{t}b(X_{k_{n}(r)}^{n})\mathrm{d}r=\int_{0}^{t}b(\varphi_{k_{n}(r)}^{n}+B_{k_{n}(r)}^{H})\mathrm{d}r,$
meanwhile, we know that our aim is to get rid of ${\mathcal{C}}^{1}$-regularity demand on $b$ via proper implement of the regularization from $B^{H}$, precisely the Gaussian density $p_{c(H)t^{2H}}$ which is infinitely many smoothing (see Section˜3.2). Also keep in mind in the current setting Girsanov Theorem is not available. What we have learned from [Ger23, BDG24] is that we can achieve this goal via freezing the exponent $\varphi_{r}$ inside the integral $\int_{0}^{t}b(\varphi_{r}+B_{r}^{H})\mathrm{d}r$ (the same for $\varphi^{n}_{k_{n}(r)}$ inside the integral $\int_{0}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$) via taking conditional expectation in the framework of SSL. That is to say, heuristically for tst-s being small enough
$\int_{s}^{t}b(\varphi_{r}+B_{r}^{H})\mathrm{d}r$ $\overset{\|\cdot\|_{L^{p}_{\omega}}}{\approx}\int_{s}^{t}{\mathbb{E}}^{s-(t-s)}b({\mathbb{E}}^{s-(t-s)}\varphi_{r}+B_{r}^{H})\mathrm{d}r,$
$\int_{s}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$ $\overset{\|\cdot\|_{L^{p}_{\omega}}}{\approx}\int_{s}^{t}{\mathbb{E}}^{s-(t-s)}b({\mathbb{E}}^{s-(t-s)}\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r,$
and this $\overset{\|\cdot\|_{L^{p}_{\omega}}}{\approx}$” is verified by SSL (see Lemma˜3.1 below) by taking
$A_{s,t}:=\int_{s}^{t}{\mathbb{E}}^{s-(t-s)}b({\mathbb{E}}^{s-(t-s)}\varphi_{r}+B_{r}^{H})\mathrm{d}r,\quad{\mathcal{A}}_{s,t}:=\int_{s}^{t}b(\varphi_{r}+B_{r}^{H})\mathrm{d}r,$
similarly for $\int_{s}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$, then together with the property of Gaussian density $p_{c(H)t^{2H}}$ of fBM ( $\mathcal{P}_{t}^{H}f:=p_{c(H)t^{2H}}\ast f$) we can further write
$A_{s,t}=\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})\mathrm{d}r.$
Now we can see that instead of dealing with $b$ directly, we have more regularity to spare in $\mathcal{P}^{H}_{t}b$ due to the smoothing effect from the convolution with $p_{c(H)t^{2H}}$.
Although the true analysis later also contains large part technical settlement, the core of the whole argument is clear. In the end we are able to adjust the idea above into the proof of the convergence rate in the following form:
$\big\|(X_{t}-X_{t}^{n})-(X_{s}-X_{s}^{n})\big\|_{L^{p}_{\omega}}=\big\|\varphi_{t}-\varphi_{t}^{n}\big\|_{L^{p}_{\omega}}$ $\leqslant N(\|\varphi-\varphi^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon}$
$=N(\|X-X^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon}$
for sufficiently small $\varepsilon$. Therefore we get (2.4) after fine dividing the interval $[0,1]$ and for each sub-interval applying what we have obtained above.
3. Auxiliary Tools
In this section, we primarily introduce our main tool which is the stochastic sewing lemma and present some properties of the fractional Brownian motions.
3.1. Stochastic Sewing Lemma
Given $M\geqslant 0$ we define $S,T]_{M}^{2}=\{(s,t)|S\leqslant s<t\leqslant T,s-M(t-s)\geqslant S\}$ and $\overline{[S,T]}_{M}^{3}=\{(s,u,t)|(s,t)\in[S,T]_{M}^{2},(u-s)\wedge(t-u)\geqslant\frac{t-s}{3}\}$.
Lemma 3.1.
[Ger23, Lemma 2.2] Let $0\leqslant S<T\leqslant 1,p\in[2,\infty),M\geqslant 0$ and let $\left(A_{s,t}\right)_{(s,t)\in[S,T]_{M}^{2}}$ be a family of random variables in $L^{p}\left(\Omega,\mathbb{R}^{d}\right)$ such that $A_{s,t}$ is $\mathcal{F}_{t}$-measurable. Suppose that for some $\varepsilon_{1},\varepsilon_{2}>0$ and $C_{1},C_{2}$ the bounds
$\big\|A_{s,t}\big\|_{L^{p}_{\omega}}\leqslant C_{1}|t-s|^{1/2+\varepsilon_{1}}$ (3.1)
and
$\big\|\mathbf{E}^{s-M(t-s)}\delta A_{s,u,t}\big\|_{L^{p}_{\omega}}\leqslant C_{2}|t-s|^{1+\varepsilon_{2}}$ (3.2)
hold for all $(s,t)\in[S,T]_{M}^{2}$ and $(s,u,t)\in\overline{[S,T]}_{M}^{3}$, where $\delta A_{s,u,t}:=A_{s,t}-A_{s,u}-A_{u,t}$. Then there exists a unique (up to modification) adapted process $\mathcal{A}:[S,T]\rightarrow L^{p}\left(\Omega,\mathbb{R}^{d}\right)$ such that $\mathcal{A}_{S}=0$ and such that for some constants $K_{1},K_{2}<\infty$, depending only on $\varepsilon_{1},\varepsilon_{2},p,d$, and $M$, the bound
$\left\|\mathcal{A}_{t}-\mathcal{A}_{s}\right\|_{L^{p}_{\omega}}\leqslant K_{1}C_{1}|t-s|^{1/2+\varepsilon_{1}}+K_{2}C_{2}|t-s|^{1+\varepsilon_{2}}$ (3.3)
holds for all $(s,t)\in[S,T]_{0}^{2}$. Moreover, if there exists any continuous process $\widetilde{\mathcal{A}}$ : $S,T]\rightarrow L^{p}\left(\Omega,\mathbb{R}^{d}\right),\varepsilon_{3}>0$, and $K_{3}<\infty$, such that $\widetilde{\mathcal{A}}_{S}=0$ and
$\big\|\widetilde{\mathcal{A}}_{t}-\widetilde{\mathcal{A}}_{s}-A_{s,t}\big\|_{L^{p}_{\omega}}\leqslant K_{3}|t-s|^{1+\varepsilon_{3}}$ (3.4)
holds for all $(s,t)\in[S,T]_{M}^{2}$, then $\widetilde{\mathcal{A}}_{t}=\mathcal{A}_{t}$ for all $S\leqslant t\leqslant T$.
3.2. Fractional Brownian motions
Here we recall the following properties concerning $B^{H}$ that have been used heavily in later analysis.
Lemma 3.2.
[…]
${\mathbb{E}}|B_{t}^{H}-{\mathbb{E}}^{s}B_{t}^{H}|^{2}=dc(H)|t-s|^{2H}\text{ and }B_{t}^{H}-{\mathbb{E}}^{s}B_{t}^{H}\text{ is independent of }\mathcal{F}_{s}.$ (3.5)
[…]
${\mathbb{E}}|B_{t}^{H}-B_{s}^{H}|\leqslant N|t-s|^{H\wedge 1}.$ (3.6)
We let $p_{t}(x)$ denote the known heat density $\frac{1}{(2\pi t)^{d/2}}e^{-\frac{|x|^{2}}{2t}}$ on ${\mathbb{R}}^{d}$ and we define ${\mathcal{P}}_{t}^{H}f(x):=(p_{c(H)t^{2H}}\ast f)(x)$, $x\in{\mathbb{R}}^{d}$. Then for any $\mathcal{F}_{s}$-measurable ${\mathbb{R}}^{d}$ valued random vector $\xi$, we have
${\mathbb{E}}^{s}f(B_{t}^{H}+\xi)={\mathcal{P}}_{t-s}^{H}f({\mathbb{E}}^{s}B_{t}^{H}+\xi).$ (3.7)
Lemma 3.3.
For $\alpha,\beta\in[0,1],f\in{\mathcal{C}}^{\alpha},t\in(0,1],$ one has the bounds, with some constant $N$ depending only on $H,\alpha,\beta,d$, accordingly
$|\mathcal{P}_{t}^{H}f(x_{1})-\mathcal{P}_{t}^{H}f(x_{2})-\mathcal{P}_{t}^{H}f(x_{3})+\mathcal{P}_{t}^{H}f(x_{4})|$
$\leqslant N\|f\|_{{\mathcal{C}}^{\alpha}}\big(t^{H(\alpha-2)}|x_{1}-x_{2}||x_{1}-x_{3}|+t^{H(\alpha-1)}|x_{1}-x_{2}-x_{3}+x_{4}|\big),\forall x_{i}\in{\mathbb{R}}^{d},i=1,\ldots,4;$ (3.8)
$\|\mathcal{P}_{t}^{H}f\|_{{\mathcal{C}}^{\beta}}\leqslant Nt^{H(\alpha-\beta)\wedge 0}\|f\|_{{\mathcal{C}}^{\alpha}};$ (3.9)
$\|(\mathcal{P}_{t}^{H}-\mathcal{P}_{s}^{H})f\|_{{\mathcal{C}}^{\beta}}\leqslant Ns^{H(\alpha-\beta)-2H\delta}|t^{2H}-s^{2H}|^{\delta}\|f\|_{{\mathcal{C}}^{\alpha}},\forall 0\leqslant s\leqslant t\leqslant 1,0<\delta\in\left[\frac{\alpha-\beta}{2},1\right].$ (3.10)
Proof.
(3.3) are directly from [Ger23, Page 2 (2.8)]. By using properties of Gaussian convolutions, heat kernel bounds and a relation of the form ${\mathcal{P}}_{t}^{H}f(x)=(p_{c(H)t^{2H}}\ast f)(x)$, we get (3.9) from [BDG21, Proposition 3.7 (i)]. For (3.10), it holds from [BDG21, Proposition 3.7 (ii)]. ∎
4. Strong convergence rate
In this part we give the proof for Theorem˜2.2.
Denote
$\varphi_{t}\coloneqq(X-B^{H})_{t}$ $=x_{0}+\int_{0}^{t}b(\varphi_{s}+B^{H}_{s})\mathrm{d}s,$
$\varphi^{n}_{t}\coloneqq(X^{n}-B^{H})_{t}$ $=x_{0}^{n}+\int_{0}^{t}b(\varphi_{k_{n}(s)}^{n}+B^{H}_{k_{n}(s)})\mathrm{d}s.$
Fix $S\leqslant s<t\leqslant T$ and $S,T]\subset[0,1]$. We write
$(X$ $-X^{n})_{t}-(X-X^{n})_{s}$
$=(\varphi-\varphi^{n})_{t}-(\varphi-\varphi^{n})_{s}$
$=\int_{s}^{t}b(B^{H}_{r}+\varphi_{r})-b(B^{H}_{r}+\varphi_{r}^{n})\mathrm{d}r+\int_{s}^{t}b(B^{H}_{r}+\varphi_{r}^{n})-b(B^{H}_{r}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$\quad+\int_{s}^{t}b(B^{H}_{r}+\varphi^{n}_{k_{n}(r)})-b(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$\eqqcolon\mathcal{E}^{b,n,1}_{s,t}+\mathcal{E}^{b,n,2}_{s,t}+\mathcal{E}^{b,n,3}_{s,t}.$
It is clear that in order to show (2.4), we need to estimate ${\mathcal{E}}^{b,n,1}_{s,t},{\mathcal{E}}^{b,n,2}_{s,t},{\mathcal{E}}^{b,n,3}_{s,t}$ individually. We distribute the estimates for each into Lemma˜4.2, Lemma˜4.3 and Lemma˜4.4 correspondingly.
Before that we first present the following auxiliary lemma for the processes $\varphi$ and $\varphi^{n}$ defined above which will be heavily used in the later proofs.
Lemma 4.1.
Assume Assumption˜2.1 holds. Then for all $t>s$ and $p\geqslant 1$ we have
$\|\varphi_{t}-{\mathbb{E}}^{s}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant N|t-s|^{1+\alpha H}\quad a.s.;$ (4.1)
$\|\varphi^{n}_{t}-{\mathbb{E}}^{s}\varphi_{t}^{n}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant N|t-s|^{1+\alpha H}\quad a.s.$ (4.2)
with some constant $N=N(p,d,\alpha,H,\|b\|_{{\mathcal{C}}^{\alpha}})$.
Proof.
Suppose (4.1) holds for some $m\geqslant 0$ in place of $1+\alpha H.$ This is certainly true for $m=0$ thanks to the fact that $b$ is bounded; we proceed now by induction on $m.$
We apply (2.2) with $X=\varphi_{t},Y=\varphi_{s}+\int_{s}^{t}b({\mathbb{E}}^{s}B^{H}_{r}+{\mathbb{E}}^{s}\varphi_{r})\mathrm{d}r$ to get
$\|\varphi_{t}-{\mathbb{E}}^{s}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant 2\left\|\varphi_{t}-\varphi_{s}-\int_{s}^{t}b({\mathbb{E}}^{s}B_{r}^{H}+{\mathbb{E}}^{s}\varphi_{r})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
$=2\left\|\int_{s}^{t}(b(B^{H}_{r}+\varphi_{r})-b({\mathbb{E}}^{s}B_{r}^{H}+{\mathbb{E}}^{s}\varphi_{r}))\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\left\|\int_{s}^{t}(|B^{H}_{r}-{\mathbb{E}}^{s}B^{H}_{r}|^{\alpha}+|\varphi_{r}-{\mathbb{E}}^{s}\varphi_{r}|^{\alpha})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(\big\||B^{H}_{r}-{\mathbb{E}}^{s}B^{H}_{r}|^{\alpha}\big\|_{L^{p}_{\omega}|\mathcal{F}_{s}}+\big\||\varphi_{r}-{\mathbb{E}}^{s}\varphi_{r}|^{\alpha}\big\|_{L^{p}_{\omega}|\mathcal{F}_{s}})\mathrm{d}r.$
Using (3.7) and the induction hypothesis, we get
$\|\varphi_{t}-{\mathbb{E}}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}\leqslant N|t-s|^{(H\alpha)\wedge(m\alpha)+1}\quad a.s.$
We note that $m_{0}=0,m_{i+1}=1+(H\alpha)\wedge(m_{i}\alpha)$ reaches $1+H\alpha$ in finitely many steps, therefore we get (4.1).
Inequality (4.2) is obtained by a similar argument. For fixed $s,$ define $s^{\prime}$ to be the smallest grid point which is bigger or equal to $s,$ that is, $s^{\prime}:=\lceil ns\rceil n^{-1}$. It is crucial to note that $\varphi_{s^{\prime}}^{n}$ is $\mathcal{F}_{s}$-measurable. We proceed by induction as before and assume that (4.2) holds for some $m\geqslant 0.$ If $s\leqslant t<s^{\prime},$ then $\varphi_{t}^{n}$ is $\mathcal{F}_{s}$-measurable. Hence $\varphi^{n}_{t}={\mathbb{E}}^{s}\varphi_{t}^{n}$ and the left-hand side of (4.2) is zero. Therefore it remains to consider the case $t\geqslant s^{\prime}.$ In this case, using again (2.2) with $X=\varphi_{t}^{n},Y=\varphi^{n}_{s^{\prime}}+\int_{s^{\prime}}^{t}b({\mathbb{E}}^{s}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s}\varphi_{k_{n}(r)}^{n})\mathrm{d}r,$ we deduce
$\|\varphi_{t}^{n}-{\mathbb{E}}^{s}\varphi_{t}^{n}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant 2\left\|\varphi_{t}^{n}-\varphi^{n}_{s^{\prime}}-\int_{s^{\prime}}^{t}b({\mathbb{E}}^{s}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s}\varphi_{k_{n}(r)}^{n})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
$=2\left\|\int_{s^{\prime}}^{t}(b(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})-b({\mathbb{E}}^{s}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s}\varphi_{k_{n}(r)}^{n})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
$\leqslant N\left\|\int_{s^{\prime}}^{t}(|B^{H}_{k_{n}(r)}-{\mathbb{E}}^{s}B^{H}_{k_{n}(r)}|^{\alpha}+|\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s}\varphi^{n}_{k_{n}(r)}|^{\alpha})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}.$
From now on we obtain (4.2) as before. ∎
Here we introduce some notations commonly used in the proofs of Lemma˜4.2, Lemma˜4.3 and Lemma˜4.4. For $(s,u,t)\in\overline{[0,1]}_{1}^{3}$, we set
$s_{1}\coloneqq s-(t-s),s_{2}\coloneqq u-(t-u),s_{3}\coloneqq s-(u-s),s_{4}\coloneqq s,s_{5}\coloneqq u,s_{6}\coloneqq t.$ (4.3)
Note by the fact $u\leqslant\frac{2}{3}s+\frac{1}{3}t$ for $(s,u,t)\in\overline{[0,1]}_{1}^{3}$, we have $s_{2}\leqslant s_{3}.$
Let us start with the estimate for $\mathcal{E}^{b,n,1}$.
Lemma 4.2.
Suppose Assumption˜2.1 holds. Then for any $p\geqslant 1$ we have
$\left\|\mathcal{E}^{b,n,1}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant N\|\varphi-\varphi^{n}\|_{C_{p}^{\frac{1}{2}}}|t-s|^{\frac{1}{2}+\varepsilon},\quad\forall(s,t)\in[0,1]_{0}^{2}$ (4.4)
with sufficiently small $\varepsilon>0$ and some constant $N=N(p,d,\alpha,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
Proof.
The idea is to apply Lemma˜3.1. Let $M=1$, $(s,t)\in[0,1]_{1}^{2}$ and
$A_{s,t}:={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}b(B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})-b(B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})\mathrm{d}r.$
We are going to verify (3.1) and (3.2). By (3.7), we see
$A_{s,t}$ $=\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})$
$\qquad-\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi^{n}_{r})\mathrm{d}r.$ (4.5)
Then by (3.9) and (2.1), we get
$\|A_{s,t}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(r-[s-(t-s)])^{-(1-\alpha)H}\|{\mathbb{E}}^{s-(t-s)}(\varphi_{r}-\varphi_{r}^{n})\|_{L^{p}_{\omega}}\mathrm{d}r$ (4.6)
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}|t-s|^{1-(1-\alpha)H}.$
Then (3.1) holds with $C_{1}=N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}$ by the fact that $1-(1-\alpha)H>\frac{1}{2}$.
Next we verify (3.2). Let $(s,u,t)\in\overline{[0,1]}_{1}^{3}$. Recall the definition of $s_{i},i=1,\dots,6$ in (4.3). We first can write
${\mathbb{E}}^{s-(t-s)}\delta A_{s,u,t}$ (4.7)
$=$ ${\mathbb{E}}^{s_{1}}{\mathbb{E}}^{s_{3}}\int_{s_{4}}^{s_{5}}b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-b(B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r})+b(B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})\mathrm{d}r$
$+{\mathbb{E}}^{s_{1}}{\mathbb{E}}^{s_{2}}\int_{s_{5}}^{s_{6}}b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-b(B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})+b(B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})\mathrm{d}r$
$=$ ${\mathbb{E}}^{s_{1}}\int_{s_{4}}^{s_{5}}\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})$
$\qquad\qquad-\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r})+\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})\mathrm{d}r$
$+{\mathbb{E}}^{s_{1}}\int_{s_{5}}^{s_{6}}\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})$
$\qquad\qquad-\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})+\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})\mathrm{d}r$
$=:$ $I_{1}+I_{2}.$
The two terms are treated in the exactly same way, so we only detail $I_{1}$. By (2.1) and applying (3.3) with
$x_{1}={\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n},x_{2}={\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r},x_{3}={\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n},x_{4}={\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r},$
we obtain
$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}|r-s_{3}|^{-(1-\alpha)H}\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}$ (4.8)
$\quad\qquad\qquad\quad+|r-s_{3}|^{-(2-\alpha)H}\big\||{\mathbb{E}}^{s_{1}}\varphi_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\cdot{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}\mathrm{d}r.$
By (4.2),
${\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|={\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}(\varphi^{n}_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})|\leqslant{\mathbb{E}}^{s_{1}}|\varphi^{n}_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\leqslant N|r-s_{1}|^{1+\alpha H}.$ (4.9)
Besides from (2.1), we get
$\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}$ $=\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}((\varphi_{r}-\varphi_{r}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}-\varphi_{r}^{n}))|\big\|_{L^{p}_{\omega}}$ (4.10)
$\leqslant\big\|{\mathbb{E}}^{s_{1}}|(\varphi_{r}-\varphi_{r}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}-\varphi_{r}^{n})|\big\|_{L^{p}_{\omega}}$
$\leqslant\|(\varphi_{r}-\varphi_{r}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}-\varphi_{r}^{n})\|_{L^{p}_{\omega}},$
meanwhile (2.3) implies
$\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}\leqslant N|r-s_{1}|^{\frac{1}{2}}[\varphi-\varphi^{n}]_{C^{\frac{1}{2}}_{p}},$ (4.11)
clearly from (2.1),
$\|{\mathbb{E}}^{s_{1}}(\varphi_{r}-\varphi_{r}^{n})\|_{L^{p}_{\omega}}\leqslant\|\varphi-\varphi^{n}\|_{C_{p}^{0}},$ (4.12)
now plugging (4.9), (4.11) and (4.12) into (4.8), we have
$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}[\varphi-\varphi^{n}]_{C_{p}^{\frac{1}{2}}}\int_{s_{4}}^{s_{5}}(r-s_{1})^{\frac{1}{2}}(r-s_{3})^{-(1-\alpha)H}\mathrm{d}r$ (4.13)
$\quad+\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}\int_{s_{4}}^{s_{5}}(r-s_{1})^{1+\alpha H}(r-s_{3})^{-(2-\alpha)H}\mathrm{d}r$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{\frac{3}{2}-(1-\alpha)H}[\varphi-\varphi^{n}]_{C_{p}^{\frac{1}{2}}}$
$\quad+\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{2+\alpha H-(2-\alpha)H}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{\frac{1}{2}}}(t-s)^{(\frac{3}{2}-(1-\alpha)H)\wedge(2+\alpha H-(2-\alpha)H)}.$
The above analysis also implies the same bound on $I_{2}$ observing $I_{1}$ and $I_{2}$ share the same structure.
Noticing Assumption˜2.1 implies $(\frac{3}{2}-(1-\alpha)H)\wedge(2+\alpha H-(2-\alpha)H)>1$, we conclude (3.2) holds with $C_{2}=N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{\frac{1}{2}}}$.
Now we claim that the process $\mathcal{A}$ in (3.3) actually is given by
${\mathcal{A}}_{t}=\int_{0}^{t}b(B^{H}_{r}+\varphi_{r})-b(B^{H}_{r}+\varphi_{r}^{n})\mathrm{d}r.$ (4.14)
To prove this, it suffices to show (3.4). By (4), we write
${\mathcal{A}}_{t}-{\mathcal{A}}_{s}-A_{s,t}$ $=\int_{s}^{t}b(B^{H}_{r}+\varphi_{r})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})\mathrm{d}r$ (4.15)
$\quad-\int_{s}^{t}b(B_{r}^{H}+\varphi_{r}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})\mathrm{d}r$
$=:II_{1}+II_{2}.$
Again we can see that $II_{2}$ can be treated similarly to $II_{1}$, so we only detail $II_{1}$. We can see
$II_{1}$ $=\int_{s}^{t}(b-\mathcal{P}^{H}_{r-s_{1}}b)(B_{r}^{H}+\varphi_{r})\mathrm{d}r+\int_{s}^{t}\mathcal{P}^{H}_{r-s_{1}}b(B_{r}^{H}+\varphi_{r})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})\mathrm{d}r.$ (4.16)
Using (3.10) with $\delta=\frac{\alpha}{2}$, $\beta=0$ and (3.9) with $\beta=\alpha$, we get
$\|II_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\int_{s}^{t}\|\mathcal{P}^{H}_{r-s_{1}}b-b\|_{{\mathcal{C}}^{0}}+\|\mathcal{P}^{H}_{r-s_{1}}b\|_{{\mathcal{C}}^{\alpha}}(\|B_{r}^{H}-{\mathbb{E}}^{s_{1}}B^{H}_{r}\|^{\alpha}_{L^{\alpha p}_{\omega}}+\|\varphi_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}\|^{\alpha}_{L_{\omega}^{\alpha p}})\mathrm{d}r$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(r-s_{1})^{\alpha H}+(|r-s_{1}|^{\alpha H}+|r-s_{1}|^{\alpha(1+\alpha H)})\mathrm{d}r$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{(1+\alpha H)\wedge(1+\alpha(1+\alpha H))},$ (4.17)
where in the second inequality we used (3.6) and (4.1). The same bound on $II_{2}$.
Therefore, (3.4) holds since $(1+\alpha H)\wedge(1+\alpha(1+\alpha H))>1$. Then the uniqueness from Lemma˜3.1 verifies the claim.
Finally, by Lemma˜3.1, the proof is complete. ∎
Let us move to estimate $\mathcal{E}^{b,n,2}$ term.
Lemma 4.3.
Suppose Assumption˜2.1 holds. Then we have for any $p\geqslant 1$ and $(s,t)\in[0,1]_{0}^{2}$
$\left\|\mathcal{E}^{b,n,2}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}|t-s|^{\frac{1}{2}+\varepsilon}$ (4.18)
where $\epsilon>0$ is sufficiently small and $N=N(p,d,\alpha,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
Proof.
Again the idea is to apply Lemma˜3.1. Let $M=1$, $(s,t)\in[0,1]_{1}^{2}$ and
$A_{s,t}:={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}b(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})-b(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$ (4.19)
Similarly to (4) and (4.6), we have
$\|A_{s,t}\|_{L^{p}_{\omega}}$ $=\|\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})$
$\qquad-\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi^{n}_{k_{n}(r)})\mathrm{d}r\|_{L^{p}_{\omega}}$ (4.20)
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\sup_{r\in[s,t]}\|\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}\|_{L_{\omega}^{p}}|t-s|^{1-(1-\alpha)H}$
$\leqslant Nn^{-1}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1-(1-\alpha)H},$ (4.21)
where in the second inequality we used $\|\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}\|_{L_{\omega}^{p}}\leqslant\|b\|_{{\mathcal{C}}^{0}}n^{-1}$.
Then, with taking $C_{1}=N\|b\|_{{\mathcal{C}}^{\alpha}}n^{-1}$, (3.1) holds by the fact that $1-(1-\alpha)H>\frac{1}{2}.$
Next we verify (3.2). Let $(s,u,t)\in\overline{[0,1]}_{1}^{3}$. Recall the definition of $s_{i},i=1,\dots,6$ in (4.3). Similarly to (4.7), we can write
${\mathbb{E}}^{s_{1}}\delta A_{s,u,t}=$ ${\mathbb{E}}^{s_{1}}\int_{s_{4}}^{s_{5}}\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
$\qquad\qquad-\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})+\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$+{\mathbb{E}}^{s_{1}}\int_{s_{5}}^{s_{6}}\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
$\qquad\qquad-\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})+\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$=$ $:I_{1}+I_{2}.$
The two terms are treated in the exactly same manner, so we only detail $I_{1}$. Similarly to (4.8), we get
$\|I_{1}\|_{L_{\omega}^{p}}\leqslant$ $N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(1-\alpha)H}\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\big\|_{L_{\omega}^{p}}$
$\qquad+(r-s_{3})^{-(2-\alpha)H}\big\||{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\cdot{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\big\|_{L_{\omega}^{p}}\mathrm{d}r.$ (4.22)
Similar to (4.10), we have
$\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\|_{L^{p}_{\omega}}\leqslant\|{\mathbb{E}}^{s_{1}}|(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})|\|_{L^{p}_{\omega}}.$ (4.23)
We note that
$\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}$ $=\int_{k_{n}(r)}^{r}b(B^{H}_{k_{n}(t)}+\varphi_{k_{n}(t)}^{n})\mathrm{d}t=(r-k_{n}(r))b(B_{k_{n}(r)}^{H}+\varphi^{n}_{k_{n}(r)})\in\mathcal{F}_{k_{n}(r)}.$
When $s_{1}\in[k_{n}(r),r],$ we have
$\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})=0;$
[…]
${\mathbb{E}}^{s_{1}}|(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})|$
$=$ $(r-k_{n}(r)){\mathbb{E}}^{s_{1}}|b(B_{k_{n}(r)}^{H}+\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}b(B_{k_{n}(r)}^{H}+\varphi_{k_{n}(r)}^{n})|$
$\leqslant$ $2(r-k_{n}(r)){\mathbb{E}}^{s_{1}}|b(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})-b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+\varphi_{s_{1}}^{n})|$
$\leqslant$ $N(r-k_{n}(r))\|b\|_{{\mathcal{C}}^{\alpha}}{\mathbb{E}}^{s_{1}}(|B_{k_{n}(r)}^{H}-{\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}|^{\alpha}+|\varphi_{k_{n}(r)}^{n}-\varphi_{s_{1}}^{n}|^{\alpha}).$
Moreover, using (3.5) and (4.2), we have
${\mathbb{E}}^{s_{1}}|(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})|$
$\leqslant$ $N(r-k_{n}(r))\|b\|_{{\mathcal{C}}^{\alpha}}(|k_{n}(r)-s_{1}|^{\alpha H}+|k_{n}(r)-s_{1}|^{\alpha})$
$\leqslant$ $\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|r-s_{1}|^{\alpha}$
where we used the fact $H>1$ in the second inequality. Plugging it into (4.23), we get
$\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|r-s_{1}|^{\alpha}.$ (4.24)
Meanwhile (4.9) and (4.2) yield
${\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\leqslant N|k_{n}(r)-s_{1}|^{1+\alpha H}\leqslant N|r-s_{1}|^{1+\alpha H},$ (4.25)
$\|{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})\|_{L_{\omega}^{p}}\leqslant N\|\varphi_{\cdot}^{n}-\varphi_{k_{n}(\cdot)}^{n}\|_{C_{p}^{0}}\leqslant N\frac{\|b\|_{C^{0}}}{n}.$ (4.26)
Applying (4.24), (4.25) and (4.26) into (4.22) gives us
$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\frac{\|b\|_{{\mathcal{C}}^{\alpha}}}{n}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(1-\alpha)H}(r-s_{1})^{\alpha}\mathrm{d}r+(r-s_{3})^{-(2-\alpha)H}(r-s_{1})^{1+\alpha H}\mathrm{d}r$
$\leqslant N\frac{\|b\|_{{\mathcal{C}}^{\alpha}}}{n}(t-s)^{(1+\alpha-(1-\alpha)H)\wedge(2+(2\alpha-2)H)}.$ (4.27)
With the same bound on $I_{2}$ and Assumption˜2.1 implying $(1+\alpha-(1-\alpha)H)\wedge(2+(2\alpha-2)H)>1$, (3.2) holds with $C_{2}=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}$.
Lastly we claim that the process $\mathcal{A}$ in (3.3) is indeed given by
$\mathcal{A}_{t}=\int_{0}^{t}b(B^{H}_{r}+\varphi_{r}^{n})-b(B^{H}_{r}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$ (4.28)
To prove this claim, it suffices to show (3.4). Similarly to (4), we have
$\mathcal{A}t-\mathcal{A}_{s}-A_{s,t}$ $=\int_{s}^{t}b(B^{H}_{r}+\varphi_{r}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi^{n}_{r})\mathrm{d}r$
$\quad-\int_{s}^{t}b(B_{r}^{H}+\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$
Then the estimates can be done similarly to that in Lemma˜4.2.
Therefore, the claim holds by the uniqueness from Lemma˜3.1. Finally, by Lemma˜3.1, we can conclude (4.18). ∎
[…]
Lemma 4.4.
Suppose Assumption˜2.1 holds. Then for any $p\geqslant 1$ and $(s,t)\in[0,1]_{0}^{2}$, we have
$\left\|\mathcal{E}^{b,n,3}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}|t-s|^{\frac{1}{2}+\varepsilon}$ (4.29)
with sufficiently small $\varepsilon>0$ and some constant $N=N(p,d,\alpha,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
Proof.
In order to apply Lemma˜3.1, this time we set
$A_{s,t}={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}b(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})-b(B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$
When $|t-s|\leqslant\frac{10}{n}$, by (3.6), we have for any $\epsilon>0$
$\|A_{s,t}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}\big\||B^{H}_{r}-B^{H}_{k_{n}(r)}|^{\alpha}\big\|_{L^{p}_{\omega}}\mathrm{d}r\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|\cdot\frac{1}{n^{\alpha}}$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\frac{1}{2}+\varepsilon}\frac{1}{n^{\alpha+\frac{1}{2}-\varepsilon}}.$
For $H>1$ and $\alpha>1-\frac{1}{2H}$ we can take such sufficiently small $\varepsilon>0$ so that $\epsilon\in(0,\alpha+\frac{1}{2H}-1)$, which implies
$\|A_{s,t}\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\frac{1}{2}+\varepsilon}.$ (4.30)
When $|t-s|>\frac{10}{n}$, we have
$A_{s,t}$ $=\int_{s_{4}}^{s_{6}}\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{k_{n}(r)-s_{1}}b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$=\int_{s_{4}}^{s_{6}}\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
$\qquad+\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{k_{n}(r)-s_{1}}b({\mathbb{E}}^{s_{1}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$=:IV_{1}+IV_{2}.$
For $IV_{1}$, by (3.9), (2.1) and (3.6), we have
$\|IV_{1}\|_{L^{p}_{\omega}}$ $\leqslant\int_{s_{4}}^{s_{6}}\|\mathcal{P}^{H}_{r-s_{1}}b\|_{{\mathcal{C}}^{1}}\|{\mathbb{E}}^{s_{1}}(B_{r}^{H}-B_{k_{n}(r)}^{H})\|_{L^{p}_{\omega}}\mathrm{d}r$
$\leqslant\frac{N\|b\|_{{\mathcal{C}}^{\alpha}}}{n}\int_{s_{4}}^{s_{6}}(r-s_{1})^{-(1-\alpha)H}\mathrm{d}r$
$\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1-(1-\alpha)H}.$ (4.31)
For $IV_{2}$, (3.10) with $\delta=1$ gives us
$\|\mathcal{P}^{H}_{t}b-\mathcal{P}_{s}^{H}b\|_{{\mathcal{C}}^{0}}\leqslant Ns^{-(2-\alpha)H}|t^{2H}-s^{2H}|\|b\|_{{\mathcal{C}}^{\alpha}},$
it implies
$|IV_{2}|$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{6}}(k_{n}(r)-s_{1})^{-(2-\alpha)H}\left((r-s_{1})^{2H}-(k_{n}(r)-s_{1})^{2H}\right)\mathrm{d}r.$
Moreover, by $k_{n}(r)-s_{1}\asymp t-s,$ and
$\big||r-s_{1}|^{2H}-|k_{n}(r)-s_{1}|^{2H}\big|\leqslant N|r-k_{n}(r)||r-s_{1}|^{2H-1}\leqslant\frac{N}{n}|t-s|^{2H-1},$
we have
$|IV_{2}|\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{6}}(t-s)^{2H-1-(2-\alpha)H}\mathrm{d}r=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{\alpha H}.$ (4.32)
Then (4.30) togehter with (4.31) and (4.32) verifies (3.1) of Lemma˜3.1 with taking $C_{1}=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}$, since $((1-(1-\alpha)H)\wedge(\alpha H\big))>\frac{1}{2}$.
Next we verify (3.2). Let $(s,u,t)\in\overline{[0,1]}_{1}^{3}$. Recall the definition of $s_{i},i=1,\dots,6$ in (4.3). Similarly to (4.7), we have
[…]
$\qquad-\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n})+\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
[…]
$\qquad-\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B^{H}_{r}+{\mathbb{E}}^{s_{2}}\varphi_{k_{n}(r)}^{n})+\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{2}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
$=:V_{1}+V_{2}.$
Again we observe that the above two terms can be treated in the exactly same way, so we only detail $V_{1}$.
Applying (3.3) with taking
$x_{1}={\mathbb{E}}^{s_{3}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)},x_{2}={\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n},$
$x_{3}={\mathbb{E}}^{s_{3}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)},x_{4}={\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n},$
we get
$\|V_{1}\|_{L^{p}_{\omega}}$
$\leqslant N\int_{s_{4}}^{s_{5}}\|b\|_{{\mathcal{C}}^{\alpha}}(r-s_{3})^{-H(2-\alpha)}\|{\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{3}}(B_{r}^{H}-B_{k_{n}(r)}^{H})|\cdot|{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)}|]\|_{L^{p}_{\omega}}\mathrm{d}r.$ (4.33)
By Cauchy-Schwarz inequality,
${\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{3}}(B_{r}^{H}-B^{H}_{k_{n}(r)})|\cdot|{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)}|]$
$\leqslant({\mathbb{E}}^{s_{1}}({\mathbb{E}}^{s_{3}}(B_{r}^{H}-B^{H}_{k_{n}(r)}))^{2})^{\frac{1}{2}}({\mathbb{E}}^{s_{1}}({\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)})^{2})^{\frac{1}{2}}.$
By Jensen inequality, we obtain
$({\mathbb{E}}^{s_{3}}(B_{r}^{H}-B^{H}_{k_{n}(r)}))^{2}\leqslant$ ${\mathbb{E}}^{s_{3}}(B_{r}^{H}-B_{k_{n}(r)}^{H})^{2},$
$({\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)})^{2}=$ $({\mathbb{E}}^{s_{3}}(\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}))^{2}\leqslant{\mathbb{E}}^{s_{3}}(\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)})^{2}.$
Therefore, by (4.2), we have
${\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{3}}(B_{r}^{H}-B^{H}_{k_{n}(r)})|\cdot|{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)}|]$
$\leqslant({\mathbb{E}}^{s_{1}}(B^{H}_{r}-B^{H}_{k_{n}(r)})^{2})^{\frac{1}{2}}\cdot({\mathbb{E}}^{s_{1}}(\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)})^{2})^{\frac{1}{2}}$
$\leqslant N({\mathbb{E}}^{s_{1}}(B^{H}_{r}-B^{H}_{k_{n}(r)})^{2})^{\frac{1}{2}}\cdot(k_{n}(r)-s_{1})^{1+\alpha H}.$
For $p\geqslant 2$, (2.1) and (3.6) imply
$\|{\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{3}}(B_{r}^{H}-B^{H}_{k_{n}(r)})||{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)}|]\|_{L^{p}_{\omega}}$ (4.34)
$\leqslant N(r-s_{1})^{1+\alpha H}\left\|\|B_{r}^{H}-B^{H}_{k_{n}(r)}\|_{L_{\omega}^{2}|\mathcal{F}_{s_{1}}}\right\|_{L^{p}_{\omega}}$
$=N(r-s_{1})^{1+\alpha H}\|{\mathbb{E}}^{s_{1}}|B_{r}^{H}-B_{k_{n}(r)}^{H}|^{2}\|^{\frac{1}{2}}_{L_{\omega}^{\frac{p}{2}}}$
$\leqslant N(r-s_{1})^{1+\alpha H}\|B_{r}^{H}-B^{H}_{k_{n}(r)}\|_{L^{p}_{\omega}}$
$\leqslant\frac{N}{n}(r-s_{1})^{1+\alpha H}.$
Therefore, plugging (4.34) into (4), we have
$\|V_{1}\|_{L^{p}_{\omega}}$ $\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(2-\alpha)H}(r-s_{1})^{1+\alpha H}\mathrm{d}r\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{2+(2\alpha-2)H}.$
The same bound holds on $V_{2}$.
Since Assumption˜2.1 implies $2+(2\alpha-2)H>1$, we can obtain that (3.2) holds with $C_{2}=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}$.
In order to show that the process $\mathcal{A}$ in (3.3) is given by
$\mathcal{A}_{t}=\int_{0}^{t}b(B^{H}_{r}+\varphi^{n}_{k_{n}(r)})-b(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r,$
it suffices to show (3.4). For doing so, we first write
$\mathcal{A}_{t}-\mathcal{A}_{s}-A_{s,t}$ $=\int_{s}^{t}b(B_{r}^{H}+\varphi^{n}_{k_{n}(r)})-\mathcal{P}_{r-s_{1}}^{H}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
$\quad-\int_{s}^{t}b(B^{H}_{k_{n}(r)}+\varphi^{n}_{k_{n}(r)})-\mathcal{P}^{H}_{(k_{n}(r)-s_{1})\vee 0}b({\mathbb{E}}^{s_{1}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r.$
Then similar to what we have done in Lemma˜4.2 (that is, (4), (4.16) and (4.17)), we can conclude the claim by the uniqueness from Lemma˜3.1.
In the end all of the conditions from Lemma˜3.1 are verified, which proves the desired result. ∎
With Lemma˜4.2, Lemma˜4.3 and Lemma˜4.4 at hand we are ready to give:
Proof of Theorem˜2.2.
By (4.4), (4.18) and (4.29), we see that
$\|(X-X^{n})_{t}-(X-X^{n})_{s}\|_{L^{p}_{\omega}}$ $\leqslant N(\|\varphi-\varphi^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon}$
$=N(\|X-X^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon},$
which implies
$[X-X^{n}]_{C_{p}^{\frac{1}{2}}[S,T]}\leqslant N(\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[S,T]}+n^{-1})(T-S)^{\varepsilon}.$
Therefore, we have
$\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[S,T]}$ $\leqslant|(X-X^{n})_{S}|+2[X-X^{n}]_{C_{p}^{\frac{1}{2}}[S,T]}$
$\leqslant|(X-X^{n})_{S}|+N(\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[S,T]}+n^{-1})(T-S)^{\varepsilon}.$
Fix $T-S=\Delta$ small enough and we obtain
$\|X-X^{n}\|_{C^{\frac{1}{2}}_{p}[S,T]}\leqslant N(|(X-X^{n})_{S}|+n^{-1}).$
Dividing $[0,1]$ into $0,\Delta],[\Delta,2\Delta],\ldots$, yields that
$\|X-X^{n}\|_{C^{\frac{1}{2}}_{p}[0,1]}\leqslant N(|x_{0}-x_{0}^{n}|+n^{-1}).$
5. Optimality
In the following we present a result concerning the optimal convergence rate of EM (1.9) for rather smooth $b\in{\mathcal{C}}^{1}$.
Theorem 5.1.
Suppose $b\in{\mathcal{C}}^{1}$ and $x_{0}^{n}=x_{0}$. Then for any $t\in[0,1]$, a.s.
$\lim_{n\rightarrow\infty}n(X_{t}-X_{t}^{n})=:c(t)$ (5.1)
exists and $c(t)$ satisfies
$c^{\prime}(t)=\nabla b(X_{t})c(t)+\frac{1}{2}\nabla b(X_{t})(b(X_{t})+(B_{t}^{H})^{\prime}),\quad c(0)=0.$
Proof.
By (1.3) and (1.9), we have
$n(X_{t}-X_{t}^{n})$ $=\int_{0}^{t}n(b\left(X_{r}\right)-b\left(X_{r}^{n}\right))+n(b\left(X_{r}^{n}\right)-b(X_{k_{n}(r)}^{n}))\mathrm{d}r$
$=\int_{0}^{t}\int_{0}^{1}\nabla b(X_{r}^{n}+\theta(X_{r}-X_{r}^{n}))n(X_{r}-X_{r}^{n})$
$\qquad\quad+\nabla b(X_{k_{n}(r)}^{n}+\theta(X_{r}^{n}-X_{k_{n}(r)}^{n}))n(X_{r}^{n}-X_{k_{n}(r)}^{n})\mathrm{d}\theta\mathrm{d}r.$
By fundamental theory of calculus we can write
$n(X_{t}-X_{t}^{n})$ $=\int_{0}^{t}\exp\left(\int_{s}^{t}\int_{0}^{1}\nabla b(X_{r}^{n}+\theta(X_{r}-X_{r}^{n}))d\theta dr\right)$
$\qquad\int_{0}^{1}\nabla b(X_{k_{n}(s)}^{n}+\theta(X_{s}^{n}-X_{k_{n}(s)}^{n}))n(X_{s}^{n}-X_{k_{n}(s)}^{n})\mathrm{d}\theta\mathrm{d}s=:c^{n}(t).$
Following from (1.9) and the fact a.s. $B^{H}\in{\mathcal{C}}^{H-\varepsilon}([0,1])$ for sufficiently small $\varepsilon>0$ by Kolmogorov continuity criteria, which means there exists a small enough $\epsilon^{\prime}>0$ so that a.s. $B_{H}\in{\mathcal{C}}^{1+\epsilon^{\prime}}$, we have a.s. (denote $\{c\}:=c-\lfloor c\rfloor$ for $c\in{\mathbb{R}}_{+}$)
$n(X_{s}^{n}-X_{k_{n}(s)}^{n})$ $=n(s-k_{n}(s))b(X_{k_{n}(s)}^{n})+n(B_{s}^{H}-B_{k_{n}(s)}^{H})$
$=\{ns\}b(X_{k_{n}(s)}^{n})+\{ns\}(B_{s}^{H})^{\prime}+o(n^{-\varepsilon^{\prime}});$
together with the condition that $\nabla b$ is continuous and bounded, moreover $X_{r}^{n}\rightarrow X_{r}$ and $X_{k_{n}(s)}^{n}\rightarrow X_{s}$ a.s. from Theorem˜2.2, dominated convergence theorem shows as $n\rightarrow\infty$, we have
$c^{n}(t)$ $-\int_{0}^{t}\exp\left(\int_{s}^{t}\nabla b(X_{r})dr\right)\nabla b(X_{s})\{ns\}(b(X_{s})+(B_{s}^{H})^{\prime})\mathrm{d}s\rightarrow 0\quad a.s.$ (5.2)
Besides, observe
$\int_{0}^{t}\exp\left(\int_{s}^{t}\nabla b(X_{r})dr\right)\nabla b(X_{s})\{ns\}(b(X_{s})+(B_{s}^{H})^{\prime})\mathrm{d}s$
$\rightarrow\frac{1}{2}\int_{0}^{t}\exp\left(\int_{s}^{t}\nabla b(X_{r})dr\right)\nabla b(X_{s})(b(X_{s})+(B_{s}^{H})^{\prime})\mathrm{d}s\quad a.s.$
since $\{n\cdot\}$ converges to $\frac{1}{2}$ weakly in $L^{2}([0,t])$ as $n\rightarrow\infty$, the above combining with (5.2) completes the proof. ∎
Remark 5.2.
Notice from Theorem˜5.1, we verify that indeed order $n$ reaches the optimal rate of EM (1.9), although for $b\in{\mathcal{C}}^{1}$ instead of assuming $b\in{\mathcal{C}}^{\alpha}$ for $\alpha<1$. Putting Theorem˜2.2 and Theorem˜5.1 together we can evidently say that order $n$ reaches the optimal (upper bound) rate of EM also for $b\in{\mathcal{C}}^{\alpha}$ simply by the embedding ${\mathcal{C}}^{1}\subset{\mathcal{C}}^{\alpha}$. However, concerning the lower bound of (1.9) for $\alpha<1$ we can not say too much so far.
Acknowledgment
CL is supported by Deutsche Forschungsgemeinschaft (DFG) - Projektnummer 563883019. KS is grateful to the financial supports by National Key R & D Program of China (No. 2022YFA1006300) and the financial supports of the NSFC (No. 12426205, No. 12271030).
We are grateful to Prof. Rongchan Zhu (Beijing Institute of Technology) for her valuable suggestions.

### v2

@@
 Abstract.
-We study a singular stochastic equation driven by a regular noise of fractional Brownian type with Hurst index $H\in(1,\infty)\setminus\mathbb{Z}$ and drift coefficient $b\in\mathcal{C}^{\alpha}$, where $\alpha>1-\tfrac{1}{2H}$.
-The strong well-posedness of this equation was first established in [Ger23], a phenomenon referred to as regularization by regular noise.
-In this note, we provide a corresponding numerical analysis.
-Specifically, we show that the Euler–Maruyama approximation $X^{n}$ converges strongly to the unique solution $X$ with rate $n^{-1}$.
-Furthermore, under the additional assumption $b\in\mathcal{C}^{1}$, we show that $n(X-X^{n})$ converges to a non-trivial limit as $n\to\infty$, thereby confirming that the rate $n^{-1}$ is in fact optimal upper bound for this scheme.
+We study a singular stochastic equation driven by regular noise of fractional Brownian type with Hurst index $H\in(1,\infty)\setminus\mathbb{Z}$ and drift coefficient $b\in\mathcal{C}^{\alpha}$, where $\alpha>1-\tfrac{1}{2H}$.
+The strong well-posedness of this equation was first established in [13], a phenomenon known as regularization by regular noise.
+In this note, we provide a numerical analysis of the equation.
+Specifically, we prove that the Euler–Maruyama approximation $X^{n}$ converges strongly to the unique solution $X$ at rate $n^{-1}$.
+Moreover, we show that $n(X-X^{n})$ converges in probability to a non-trivial limit as $n\to\infty$, which confirms that the rate $n^{-1}$ is optimal for this scheme.
+In this sense, this provides a first-order numerical method for equations with non-Lipschitz drift while still achieving the rate $n^{-1}$.
 Mathematics Subject Classification (2020): Primary 60H35, 65C30, 60H10; Secondary 60H50, 60L90.
@@
 1. Introduction
-It is known that when $H\in(0,1)$, a $d$-dimensional fractional Brwonian motion (fBM) $B^{H}$ can be defined via the Mandelbrot - van Ness representation [MVN68]:
+It is known that when $H\in(0,1)$, a $d$-dimensional fractional Brownian motion (fBM) $B^{H}$ can be defined via the Mandelbrot–van Ness representation [26]:
 $B_{t}^{H}:=\int_{-\infty}^{0}(|t-s|^{H-\frac{1}{2}}-|s|^{H-\frac{1}{2}})\mathrm{d}W_{s}+\int_{0}^{t}|t-s|^{H-\frac{1}{2}}\mathrm{d}W_{s},\quad t\geqslant 0,$ (1.1)
@@
 $B_{t}^{H}:=\int_{0\leqslant s_{1}\leqslant\ldots\leqslant s_{\lfloor H\rfloor}\leqslant t}B^{H-\lfloor H\rfloor}_{s_{1}}\mathrm{d}{s_{1}}\ldots\mathrm{d}s_{\lfloor H\rfloor}.$ (1.2)
-Due to the possible multiple times of integral in (1.2) we can see clearly that the path of $B^{H}$ for $H>1$ is regular, at least $C^{1}$.
-This fact yields one of the appealing points of [Ger23] which shows the strong well-posedness of following equation with singular $b$:
+Due to the possible multiple integrals in (1.2), we can see clearly that the paths of $B^{H}$ for $H>1$ are regular, at least $C^{1}$.
+This fact leads to one of the appealing results of [13], which shows the strong well-posedness of the following equation with singular $b$:
 $\mathrm{d}X_{t}=b(X_{t})\mathrm{d}t+\mathrm{d}B_{t}^{H},\quad X_{0}=x_{0}\in{\mathbb{R}}^{d},$ (1.3)
@@
 (1.8)
-which shares the similar framework from [CdRHM22, HZZZ24, HLL24, Lin24] but with non-Markvin noise when $H\neq k+\frac{1}{2}$ for any $k\in{\mathbb{N}}$.
-In this note our motivation is to provide a numerical result of it by considering its Euler-Maruyama (EM) scheme
+which shares a similar framework with [6, 16, 15, 25], but with non-Markovian noise when $H\neq k+\frac{1}{2}$ for any $k\in{\mathbb{N}}$.
+Our motivation here is to provide a numerical result for this equation by considering its Euler–Maruyama (EM) scheme
 $\mathrm{d}X_{t}^{n}=b(X_{k_{n}(t)}^{n})\mathrm{d}t+\mathrm{d}B_{t}^{H},\qquad X_{0}^{n}=x_{0}^{n}\in{\mathbb{R}}^{d}$ (1.9)
 with $k_{n}(t)\coloneqq\frac{\lfloor nt\rfloor}{n}$.
-When $H\in(0,1)$, [BDG21] has shown the strong convergence rate $n^{-(\frac{1}{2}+\alpha H)\wedge 1+\epsilon}$ of the scheme (1.9) to (1.3), inside which the Girsanov Theroem and Stochastic Sewing Lemma (SSL) [Lê20] play one of the most important roles.
-Precisely the methodology inside [BDG21] can be roughly summarized as (take $x_{0}=x_{0}^{n}$)
+When $H\in(0,1)$, [4] has shown the strong convergence rate $n^{-(\frac{1}{2}+\alpha H)\wedge 1+\epsilon}$ of the scheme (1.9) to (1.3), in which the Girsanov Theorem and the Stochastic Sewing Lemma (SSL) [22] play crucial roles.
+More precisely, the methodology in [4] can be roughly summarized as follows (taking $x_{0}=x_{0}^{n}$):
 $\|\sup_{t\in[0,1]}|X_{t}-X^{n}_{t}|\|_{L^{p}_{\omega}}\overset{\text{Girsanov}}{\lesssim}\big\|\int_{0}^{1}b(B_{s}^{H})-b(B_{k_{n}(s)}^{H})\mathrm{d}r\big\|_{L^{p}_{\omega}}\overset{\text{SSL}}{\lesssim}n^{-(\frac{1}{2}+\alpha H)\wedge 1+\epsilon}.$
-When $H>1$, on one hand, as indicated in [Ger23, GG25] the Girsanov Theorem becomes less helpful; meanwhile PDE tools clearly would not work.
-Therefore the challenge for showing the convergence of (1.9) to (1.3) lies in two folds comparing with known results: absence of Girsanov and non-Markovianity.
-Alternatively, although [Ger23] studies the well-posedness only, actually it already hints the possible way to get rid of Girsanov; concerning numerical approximation [BDG24] provides the similar evidence, however it considers the singular SDE driven by an $\alpha$ stable process which is a Markov process.
-Nevertheless we are able to show convergence rate $n^{-1}$ which is comparable with [BDG21] since $H>1$ here implies $\alpha>1-\frac{1}{2H}>\frac{1}{2}$.
-For such obtained convergence rate, naturally we wonder how far it is away from being optimal.
-Such question on the optimality of EM scheme has been addressed and studied for instance in [MGY23, EMGY25, DGL23a] concerning Brownian noise, [PSS24] for Lévy processes with jumps and [Neu06, NN07, LLT25, HLN16] for fBM with $H<1$.
-In particular [Neu06] confirms for $H\in(\frac{1}{2},1)$, order $n^{H}$ reaches its optimality of EM scheme (1.9) for $b\in{\mathcal{C}}^{2}$ (twice differentiable) with possible linear growth.
-Here we complete this result with rate $n^{-1}$ for $H>1$ and less regular $b$ (actually only ${\mathcal{C}}^{1}$).
-The idea of the proof is straightforward: we show that for $b\in{\mathcal{C}}^{1}$, the following approximation reaches a possibly non-zero limit:
-$n(X_{t}-X^{n}_{t})\overset{n\rightarrow\infty}{\rightarrow}c(t)\neq 0,\quad a.s.$
-which indicates that for (1.9) the best convergence rate we can expect should be not faster than order $n$. Evidently it also verifies that the rate we obtained actually is optimal.
-Lastly, let us mention a few related works that share a similar interest and spirit in the study of numerical approximations.
-For equations of the type (1.3) with singular drift, convergence results have been established in [NS21, DG20, JM21, BW25] for additive Brownian motion, in [DGL23b, BDG21, GLL25, LL25] for multiplicative Brownian noise, in [BDG24, BWWZ25] for Lévy processes with jumps, and in [BDG21, GHR25] for fractional Brownian motion.
-A slightly different notion of singularity-namely, piecewise Lipschitz coefficients-has also been investigated, with convergence results obtained in works such as [LS16, MGY20, MGRY25, PSS24].
-We must acknowledge that this is only a brief selection from a vast body of literature, and we encourage readers to explore the cited works for further details and insights.
+When $H>1$, on the one hand, as indicated in [13, 11], the Girsanov Theorem becomes less helpful; meanwhile, PDE tools clearly do not apply.
+Therefore, the challenge in showing the convergence of (1.9) to (1.3) lies in two aspects compared with known results: the absence of Girsanov’s theorem and the non-Markovian nature of the noise.
+Alternatively, although [13] studies only well-posedness, it already hints at a possible way to avoid using Girsanov’s theorem; concerning numerical approximation, [3] provides similar evidence, but it considers singular SDEs driven by an $\alpha$-stable process, which is Markovian.
+Nevertheless, we are able to show a convergence rate of $n^{-1}$, which is comparable with [4], since $H>1$ here implies $\alpha>1-\frac{1}{2H}>\frac{1}{2}$.
+Given this convergence rate, it is natural to ask how far it is from being optimal.
+Such questions on the optimality of the EM scheme have been addressed, for instance, in [31, 10, 7] for Brownian noise, [35] for Lévy processes with jumps, and [34, 32, 24, 19] for fBM with $H<1$.
+In particular, [34] confirms that for $H\in(\frac{1}{2},1)$, the order $n^{H}$ is optimal for the EM scheme (1.9) when $b\in{\mathcal{C}}^{2}$ (twice differentiable) with possible linear growth.
+Here, we complete this result with rate $n^{-1}$ for $H>1$ and less regular $b$ (in fact, only ${\mathcal{C}}^{1}$).
+The idea of the proof is straightforward: we show that for $b\in{\mathcal{C}}^{\alpha}$, the following approximation converges in probability to a possibly non-zero limit:
+$n(X_{t}-X^{n}_{t})\overset{n\rightarrow\infty}{\rightarrow}c(t)\neq 0,$
+which indicates that for (1.9), the best convergence rate one can expect is no faster than order $n^{-1}$.
+Consequently, this also verifies that the rate we obtain is indeed optimal.
+Evidently, (1.9) provides a stochastic numerical method for simulating equations with singular $b$ while still converging with rate $n^{-1}$.
+Lastly, we mention a few related works that share a similar interest and spirit in the study of numerical approximations.
+For equations of the type (1.3) with singular drift, convergence results have been established in [33, 9, 20, 2] for additive Brownian motion, in [8, 4, 12, 21] for multiplicative Brownian noise, in [3, 1] for Lévy processes with jumps, and in [4, 14] for fractional Brownian motion.
+A slightly different notion of singularity-namely, piecewise Lipschitz coefficients-has also been investigated, with convergence results obtained in works such as [23, 30, 29, 35].
+We emphasize that this is only a brief selection from a vast body of literature, and we encourage readers to consult the cited works for further details and insights.
 Organization of the paper
@@
 Section˜3 collects all of the crucial tools and properties of fBM.
-We present the central analysis and proof of the main result in Section˜4.
-In the end we discuss optimality in Section˜5.
+We present the central analysis and the proof of the main convergence result in Section˜4.
+Finally, we show optimality in Section˜5.
+Appendix˜A contains several auxiliary proofs.
 2. Preliminaries and main results
@@
 If $k=(0,\ldots,0)$, we use convention $\partial^{k}f=f$.
-We denote by ${C}^{\infty}_{0}$ (${C}_{p}^{\infty}$, resp.) the set of all continuously infinitely differentiable functions that, along with all of their partial derivatives, are compactly supported (of polynomial growth, resp.).
-For $\alpha\in(0,1]$, we set $\mathcal{C}^{\alpha}({\mathbb{R}}^{d})$ to be the space of continuous functions such that
+We denote by ${{\mathcal{C}}}^{\infty}_{0}$ (${{\mathcal{C}}}_{p}^{\infty}$, resp.) the set of all continuously infinitely differentiable functions that, along with all of their partial derivatives, are compactly supported (of polynomial growth, resp.).
+For $\alpha\in(0,1)$, we set $\mathcal{C}^{\alpha}({\mathbb{R}}^{d})$ to be the space of continuous functions such that
 $\|f\|_{\mathcal{C}^{\alpha}}\coloneqq[f]_{\mathcal{C}^{\alpha}}+\sup_{x\in{\mathbb{R}}^{d}}|f(x)|\coloneqq\sup_{x,y\in{\mathbb{R}}^{d},x\neq y}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}+\sup_{x\in{\mathbb{R}}^{d}}|f(x)|<\infty.$
@@
 To be noticed that the functions in $\mathcal{C}^{0}$ do not need to be continuous.
+Let $n\in{\mathbb{N}}$ and $\alpha,\beta\in(0,1)$ such that $\alpha+\beta>1$.
+Then for $f\in{\mathcal{C}}^{\alpha}([0,T],{\mathbb{R}}^{n})$, $g\in{\mathcal{C}}^{\beta}([0,T],{\mathbb{R}}^{n})$, the Young integral $h_{t}=\int_{0}^{t}f_{t}dg_{t}$ is well-defined as the limit as $m\rightarrow\infty$ of the Riemann sums
+$\sum_{i=0}^{m}f_{t_{i}^{m}}(g_{t_{i+1}^{m}\wedge t}-g_{t_{i}^{m}\wedge t})$
+where $(t_{i}^{m})_{0\leqslant m}$ is any partition sequence of $[0,t]$.
+And the Young integral satisfies the estimates: there exists the constant $C$ depends only on $\alpha,\beta$ so that for all $0\leqslant s\leqslant t\leqslant T,$
+$|h_{t}-h_{s}-f_{s}(g_{t}-g_{s})|\leqslant C|t-s|^{\alpha+\beta}[f]_{{\mathcal{C}}^{\alpha}([s,t])}[g]_{{\mathcal{C}}^{\beta}([s,t])},$
+which also yields the following
+$[h]_{{\mathcal{C}}^{\beta}([s,t])}\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}([s,t])}[g]_{{\mathcal{C}}^{\beta}([s,t])}.$ (2.1)
 In the following we denote the conditional expectation w.r.t.
@@
 2.2. Main results
-Our main assumption and result can be stated as follows.
+By scaling we can always take $t\in[0,1]$ without lost of generality. Our main assumption and results are stated as follows.
 Assumption 2.1.
 […]
-Notice that following [Ger23], under the assumption above, there exists a unique strong solution to (1.3). Here is our numerical approximation result for it.
+Notice that following [13], under the assumption above, there exists a unique strong solution to (1.3). Here is our numerical approximation result for it.
 Theorem 2.2.
-Let $(X_{t})_{t\in[0,1]},(X_{t}^{n})_{t\in[0,1]}$ be the solutions to (1.3) and (1.9) accordingly. Suppose Assumption˜2.1 holds. Then for every $p>1$, we have
-$\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[0,1]}\leqslant N|x_{0}-x_{0}^{n}|+Nn^{-1},$ (2.4)
-where $N=N(p,d,\alpha,H,\|b\|_{{\mathcal{C}}^{\alpha}})$.
+Let $(X_{t})_{t\in[0,1]},(X_{t}^{n})_{t\in[0,1]}$ be the solutions to (1.3) and (1.9) accordingly. Suppose Assumption˜2.1 holds. Then for every $p\geqslant 1$ and $\gamma<1+(\alpha-1)H$, we have
+$\|X-X^{n}\|_{C_{p}^{\gamma}[0,1]}\leqslant C|x_{0}-x_{0}^{n}|+Cn^{-1},$ (2.5)
+where $C=C(p,d,\alpha,H,\|b\|_{{\mathcal{C}}^{\alpha}})$.
 Remark 2.3.
 As an easy application of Kolmogorov continuity criteria we can also conclude from (2.5) the following:
-$\big\|\sup_{t\in[0,1]}|X_{t}-X_{t}^{n}|\big\|_{L_{\omega}^{p}}\leqslant N|x_{0}-x_{0}^{n}|+Nn^{-1}.$
-Idea of the analysis. Its detailed proof will be given in Section˜4. Here we only outline the rough idea.
-Observe that for any $p\geqslant 1$, for $0\leqslant s\leqslant t\leqslant 1$
-$\big\|(X_{t}-X_{t}^{n})-(X_{s}-X_{s}^{n})|\big\|_{L^{p}_{\omega}}=\big\|(\varphi_{t}-\varphi_{t}^{n})-(\varphi_{s}-\varphi_{s}^{n})\big\|_{L^{p}_{\omega}},$
+$\big\|\sup_{t\in[0,1]}|X_{t}-X_{t}^{n}|\big\|_{L_{\omega}^{p}}\leqslant C|x_{0}-x_{0}^{n}|+Cn^{-1}.$
+The following theorem settles the question on optimality.
+Theorem 2.4.
+Let $(X_{t})_{t\in[0,1]},(X_{t}^{n})_{t\in[0,1]}$ be the solutions to (1.3) and (1.9) accordingly and $x_{0}=x_{0}^{n}$.
+Suppose Assumption˜2.1 holds.
+Then for every $f\in{\mathcal{C}}^{\alpha}({\mathbb{R}}^{d})$ there exists a $({\mathcal{F}}_{t})_{t\in[0,1]}$-adapted process $({\mathcal{H}}f)^{X}_{t\in[0,1]}\in C_{p}^{1+(\alpha-1)H-}[0,1]$ for every $p\geqslant 2$ such that bounded linearly
+${\mathcal{C}}^{\alpha}\ni f\mapsto({\mathcal{H}}f)^{X}\in C_{p}^{1+(\alpha-1)H-}[0,1]$
+and if $g\in{\mathcal{C}}^{1}({\mathbb{R}}^{d})$, with probability one
+$({\mathcal{H}}g)_{t}^{X}=\int_{0}^{t}\nabla g(X_{s})\mathrm{d}s,\quad t\in[0,1];$
+moreover, we have for any $t\in[0,1]$, in probability
+$\lim_{n\rightarrow\infty}n(X_{t}-X_{t}^{n})=:c(t)$ (2.6)
+exists and $c\in{\mathcal{C}}^{1+(\alpha-1)H-}([0,1])$ satisfies
+$c(t)=\int_{0}^{t}c(s)\mathrm{d}({\mathcal{H}}b)^{X}_{s}+\frac{1}{2}(b(X_{t})-b(x_{0})),\quad t\in[0,1],$ (2.7)
+where the integral inside (2.7) is defined in the sense of Young.
+Idea of the analysis. The detailed proofs of Theorem˜2.2 and Theorem˜2.4 will be given in Section˜4 and Section˜5, respectively. Here we only outline the main ideas.
+  1. (1)
+    (Convergence) Observe that for any $p\geqslant 1$, for $0\leqslant s\leqslant t\leqslant 1$,
+    $\big\|(X_{t}-X_{t}^{n})-(X_{s}-X_{s}^{n})\big\|_{L^{p}_{\omega}}=\big\|(\varphi_{t}-\varphi_{t}^{n})-(\varphi_{s}-\varphi_{s}^{n})\big\|_{L^{p}_{\omega}},$
     where
@@
     $\varphi_{t}^{n}:=X_{t}^{n}-B_{t}^{H}=\int_{0}^{t}b(X_{k_{n}(r)}^{n})\mathrm{d}r=\int_{0}^{t}b(\varphi_{k_{n}(r)}^{n}+B_{k_{n}(r)}^{H})\mathrm{d}r,$
-meanwhile, we know that our aim is to get rid of ${\mathcal{C}}^{1}$-regularity demand on $b$ via proper implement of the regularization from $B^{H}$, precisely the Gaussian density $p_{c(H)t^{2H}}$ which is infinitely many smoothing (see Section˜3.2).
-Also keep in mind in the current setting Girsanov Theorem is not available.
-What we have learned from [Ger23, BDG24] is that we can achieve this goal via freezing the exponent $\varphi_{r}$ inside the integral $\int_{0}^{t}b(\varphi_{r}+B_{r}^{H})\mathrm{d}r$ (the same for $\varphi^{n}_{k_{n}(r)}$ inside the integral $\int_{0}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$) via taking conditional expectation in the framework of SSL.
-That is to say, heuristically for tst-s being small enough
+    meanwhile, we know that our aim is to remove the ${\mathcal{C}}^{1}$ regularity requirement on $b$ via proper use of the regularization coming from $B^{H}$, namely the Gaussian density $p_{c(H)t^{2H}}$, which is infinitely smoothing (see Section˜3.2).
+Also keep in mind that in the current setting the Girsanov Theorem is not available.
+What we have learned from [13, 3] is that we can achieve this goal via freezing the exponent $\varphi_{r}$ inside the integral $\int_{0}^{t}b(\varphi_{r}+B_{r}^{H})\mathrm{d}r$ (and similarly for $\varphi^{n}_{k_{n}(r)}$ inside the integral $\int_{0}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$) by taking conditional expectation in the framework of the SSL.
+That is to say, heuristically, for tst-s small enough,
     $\int_{s}^{t}b(\varphi_{r}+B_{r}^{H})\mathrm{d}r$ $\overset{\|\cdot\|_{L^{p}_{\omega}}}{\approx}\int_{s}^{t}{\mathbb{E}}^{s-(t-s)}b({\mathbb{E}}^{s-(t-s)}\varphi_{r}+B_{r}^{H})\mathrm{d}r,$
     $\int_{s}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$ $\overset{\|\cdot\|_{L^{p}_{\omega}}}{\approx}\int_{s}^{t}{\mathbb{E}}^{s-(t-s)}b({\mathbb{E}}^{s-(t-s)}\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r,$
-and this $\overset{\|\cdot\|_{L^{p}_{\omega}}}{\approx}$” is verified by SSL (see Lemma˜3.1 below) by taking
+    and this $\overset{\|\cdot\|_{L^{p}_{\omega}}}{\approx}$” is justified by the SSL (see Lemma˜3.1 below) by taking
     $A_{s,t}:=\int_{s}^{t}{\mathbb{E}}^{s-(t-s)}b({\mathbb{E}}^{s-(t-s)}\varphi_{r}+B_{r}^{H})\mathrm{d}r,\quad{\mathcal{A}}_{s,t}:=\int_{s}^{t}b(\varphi_{r}+B_{r}^{H})\mathrm{d}r,$
-similarly for $\int_{s}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$, then together with the property of Gaussian density $p_{c(H)t^{2H}}$ of fBM ( $\mathcal{P}_{t}^{H}f:=p_{c(H)t^{2H}}\ast f$) we can further write
+    and similarly for $\int_{s}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$.
+Then, together with the property of the Gaussian density $p_{c(H)t^{2H}}$ of fBM ( $\mathcal{P}_{t}^{H}f:=p_{c(H)t^{2H}}\ast f$), we can further write
     $A_{s,t}=\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})\mathrm{d}r.$
-Now we can see that instead of dealing with $b$ directly, we have more regularity to spare in $\mathcal{P}^{H}_{t}b$ due to the smoothing effect from the convolution with $p_{c(H)t^{2H}}$.
-Although the true analysis later also contains large part technical settlement, the core of the whole argument is clear.
-In the end we are able to adjust the idea above into the proof of the convergence rate in the following form:
-$\big\|(X_{t}-X_{t}^{n})-(X_{s}-X_{s}^{n})\big\|_{L^{p}_{\omega}}=\big\|\varphi_{t}-\varphi_{t}^{n}\big\|_{L^{p}_{\omega}}$ $\leqslant N(\|\varphi-\varphi^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon}$
-$=N(\|X-X^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon}$
-for sufficiently small $\varepsilon$. Therefore we get (2.4) after fine dividing the interval $[0,1]$ and for each sub-interval applying what we have obtained above.
+    Now we can see that instead of dealing with $b$ directly, we gain additional regularity in $\mathcal{P}^{H}_{t}b$ due to the smoothing effect of the convolution with $p_{c(H)t^{2H}}$.
+    Although the full analysis later also contains many technical details, the core of the argument is clear.
+In the end, we are able to turn the idea above into a proof of the convergence rate in the following form:
+    $\big\|(X_{t}-X_{t}^{n})-(X_{s}-X_{s}^{n})\big\|_{L^{p}_{\omega}}=\big\|\varphi_{t}-\varphi_{t}^{n}\big\|_{L^{p}_{\omega}}$ $\leqslant C(\|\varphi-\varphi^{n}\|_{C^{\gamma}_{p}[s,t]}+n^{-1})|t-s|^{\gamma+\varepsilon}$
+    $=C(\|X-X^{n}\|_{C^{\gamma}_{p}[s,t]}+n^{-1})|t-s|^{\gamma+\varepsilon}$
+    for sufficiently small $\varepsilon$. Therefore, we obtain (2.5) after finely dividing the interval $[0,1]$ and applying the above estimate on each sub-interval.
+  2. (2)
+    (Optimality) Our idea for verifying optimality is to establish a limit theorem for the asymptotic error distribution of $X^{n}$ and $X$, that is, to show that $n(X^{n}-X)$ converges to a nontrivial limit, indicating that order $n$ is the optimal convergence rate of $X^{n}$ from (1.9) to $X$.
+    For this, morally, if the coefficient $b$ is smooth, we can show that there exists a nonzero process $c(t)$ such that $c(t)=\lim_{n\rightarrow\infty}n(X_{t}-X_{t}^{n})$ in probability and it satisfies a random ODE
+    $c(t)=\int_{0}^{t}\nabla b(X_{s})c(s)ds+\frac{1}{2}(b(X_{t})-b(x_{0})).$ (2.8)
+    At this point, we already see the difficulty in our setting: $b\in{\mathcal{C}}^{\alpha}$ with $\alpha\in(1-\frac{1}{2H},1)$ implies that $\nabla b(X)$ is ill-defined in the above equation.
+    However, if we reformulate (2.8) equivalently as follows: denote $H_{t}:=\int_{0}^{t}\nabla b(X_{s})\mathrm{d}s$,
+    $c(t)=\int_{0}^{t}c(s)\mathrm{d}H_{s}+\frac{1}{2}(b(X_{t})-b(x_{0})),$ (2.9)
+    then, due to the regularization effect, we can show that a.s.
+$H$ is ${\mathcal{C}}^{1+(\alpha-1)H-}$-Hölder continuous (see Lemma˜5.1), which implies that the term $\int_{0}^{t}c(s)\mathrm{d}H_{s}$ can be defined in the sense of the Young integral.
+Therefore, we can claim that (2.8) is well defined, given that (2.9) is a linear equation.
+In this way, we also show that $n(X^{n}-X)$ indeed converges to a nontrivial limit and thus that order $n$ is optimal.
+    Let us also point out that the general idea of this part is close to that in [7].
+However, here we are dealing with additive fBM noise, whereas [7] considers central limit theorem type results that heavily rely on the multiplicative structure of the noise; therefore, the detailed and essential analysis deviates.
+In the end, unlike the weakly convergence result from [7], we are able to show that $n(X^{n}_{t}-X_{t})$ converges to $c(t)$ in probability.
 3. Auxiliary Tools
@@
 Lemma 3.1.
-[Ger23, Lemma 2.2] Let $0\leqslant S<T\leqslant 1,p\in[2,\infty),M\geqslant 0$ and let $\left(A_{s,t}\right)_{(s,t)\in[S,T]_{M}^{2}}$ be a family of random variables in $L^{p}\left(\Omega,\mathbb{R}^{d}\right)$ such that $A_{s,t}$ is $\mathcal{F}_{t}$-measurable.
+[13, Lemma 2.2] Let $0\leqslant S<T\leqslant 1,p\in[2,\infty),M\geqslant 0$ and let $\left(A_{s,t}\right)_{(s,t)\in[S,T]_{M}^{2}}$ be a family of random variables in $L^{p}\left(\Omega,\mathbb{R}^{d}\right)$ such that $A_{s,t}$ is $\mathcal{F}_{t}$-measurable.
 Suppose that for some $\varepsilon_{1},\varepsilon_{2}>0$ and $C_{1},C_{2}$ the bounds
@@
 3.2. Fractional Brownian motions
-Here we recall the following properties concerning $B^{H}$ that have been used heavily in later analysis.
+We recall the following properties concerning $B^{H}$ that have been used heavily in later analysis.
 Lemma 3.2.
@@
 […]
-${\mathbb{E}}|B_{t}^{H}-B_{s}^{H}|\leqslant N|t-s|^{H\wedge 1}.$ (3.6)
+${\mathbb{E}}|B_{t}^{H}-B_{s}^{H}|\leqslant C|t-s|^{H\wedge 1}.$ (3.6)
 We let $p_{t}(x)$ denote the known heat density $\frac{1}{(2\pi t)^{d/2}}e^{-\frac{|x|^{2}}{2t}}$ on ${\mathbb{R}}^{d}$ and we define ${\mathcal{P}}_{t}^{H}f(x):=(p_{c(H)t^{2H}}\ast f)(x)$, $x\in{\mathbb{R}}^{d}$.
@@
 Lemma 3.3.
-For $\alpha,\beta\in[0,1],f\in{\mathcal{C}}^{\alpha},t\in(0,1],$ one has the bounds, with some constant $N$ depending only on $H,\alpha,\beta,d$, accordingly
+For $\alpha,\beta\in[0,1],f\in{\mathcal{C}}^{\alpha},t\in(0,1],$ one has the bounds, with some constant $C$ depending only on $H,\alpha,\beta,d$, accordingly
 $|\mathcal{P}_{t}^{H}f(x_{1})-\mathcal{P}_{t}^{H}f(x_{2})-\mathcal{P}_{t}^{H}f(x_{3})+\mathcal{P}_{t}^{H}f(x_{4})|$
-$\leqslant N\|f\|_{{\mathcal{C}}^{\alpha}}\big(t^{H(\alpha-2)}|x_{1}-x_{2}||x_{1}-x_{3}|+t^{H(\alpha-1)}|x_{1}-x_{2}-x_{3}+x_{4}|\big),\forall x_{i}\in{\mathbb{R}}^{d},i=1,\ldots,4;$ (3.8)
-$\|\mathcal{P}_{t}^{H}f\|_{{\mathcal{C}}^{\beta}}\leqslant Nt^{H(\alpha-\beta)\wedge 0}\|f\|_{{\mathcal{C}}^{\alpha}};$ (3.9)
-$\|(\mathcal{P}_{t}^{H}-\mathcal{P}_{s}^{H})f\|_{{\mathcal{C}}^{\beta}}\leqslant Ns^{H(\alpha-\beta)-2H\delta}|t^{2H}-s^{2H}|^{\delta}\|f\|_{{\mathcal{C}}^{\alpha}},\forall 0\leqslant s\leqslant t\leqslant 1,0<\delta\in\left[\frac{\alpha-\beta}{2},1\right].$ (3.10)
+$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\big(t^{H(\alpha-2)}|x_{1}-x_{2}||x_{1}-x_{3}|+t^{H(\alpha-1)}|x_{1}-x_{2}-x_{3}+x_{4}|\big),\forall x_{i}\in{\mathbb{R}}^{d},i=1,\ldots,4;$ (3.8)
+$\|\mathcal{P}_{t}^{H}f\|_{{\mathcal{C}}^{\beta}}\leqslant Ct^{H(\alpha-\beta)\wedge 0}\|f\|_{{\mathcal{C}}^{\alpha}};$ (3.9)
+$\|(\mathcal{P}_{t}^{H}-\mathcal{P}_{s}^{H})f\|_{{\mathcal{C}}^{\beta}}\leqslant Cs^{H(\alpha-\beta)-2H\delta}|t^{2H}-s^{2H}|^{\delta}\|f\|_{{\mathcal{C}}^{\alpha}},\forall 0\leqslant s\leqslant t\leqslant 1,0<\delta\in\left[\frac{\alpha-\beta}{2},1\right].$ (3.10)
 Proof.
-(3.3) are directly from [Ger23, Page 2 (2.8)].
-By using properties of Gaussian convolutions, heat kernel bounds and a relation of the form ${\mathcal{P}}_{t}^{H}f(x)=(p_{c(H)t^{2H}}\ast f)(x)$, we get (3.9) from [BDG21, Proposition 3.7 (i)].
-For (3.10), it holds from [BDG21, Proposition 3.7 (ii)].
+(3.3) are directly from [13, Page 2 (2.8)].
+By using properties of Gaussian convolutions, heat kernel bounds and a relation of the form ${\mathcal{P}}_{t}^{H}f(x)=(p_{c(H)t^{2H}}\ast f)(x)$, we get (3.9) from [4, Proposition 3.7 (i)].
+For (3.10), it holds from [4, Proposition 3.7 (ii)].
 ∎
@@
 Lemma 4.1.
-Assume Assumption˜2.1 holds. Then for all $t>s$ and $p\geqslant 1$ we have
-$\|\varphi_{t}-{\mathbb{E}}^{s}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant N|t-s|^{1+\alpha H}\quad a.s.;$ (4.1)
-$\|\varphi^{n}_{t}-{\mathbb{E}}^{s}\varphi_{t}^{n}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant N|t-s|^{1+\alpha H}\quad a.s.$ (4.2)
-with some constant $N=N(p,d,\alpha,H,\|b\|_{{\mathcal{C}}^{\alpha}})$.
+Assume Assumption˜2.1 holds. Then for all $t>s$ and $p\geqslant 1$ we have a.s.
+$\|\varphi_{t}-{\mathbb{E}}^{s}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant C\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1+\alpha H};$ (4.1)
+$\|\varphi^{n}_{t}-{\mathbb{E}}^{s}\varphi_{t}^{n}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant C\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1+\alpha H}$ (4.2)
+with some constant $C=C(p,d,\alpha,H)$.
 Proof.
-Suppose (4.1) holds for some $m\geqslant 0$ in place of $1+\alpha H.$ This is certainly true for $m=0$ thanks to the fact that $b$ is bounded; we proceed now by induction on $m.$
-We apply (2.2) with $X=\varphi_{t},Y=\varphi_{s}+\int_{s}^{t}b({\mathbb{E}}^{s}B^{H}_{r}+{\mathbb{E}}^{s}\varphi_{r})\mathrm{d}r$ to get
-$\|\varphi_{t}-{\mathbb{E}}^{s}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant 2\left\|\varphi_{t}-\varphi_{s}-\int_{s}^{t}b({\mathbb{E}}^{s}B_{r}^{H}+{\mathbb{E}}^{s}\varphi_{r})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
-$=2\left\|\int_{s}^{t}(b(B^{H}_{r}+\varphi_{r})-b({\mathbb{E}}^{s}B_{r}^{H}+{\mathbb{E}}^{s}\varphi_{r}))\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
-$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\left\|\int_{s}^{t}(|B^{H}_{r}-{\mathbb{E}}^{s}B^{H}_{r}|^{\alpha}+|\varphi_{r}-{\mathbb{E}}^{s}\varphi_{r}|^{\alpha})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
-$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(\big\||B^{H}_{r}-{\mathbb{E}}^{s}B^{H}_{r}|^{\alpha}\big\|_{L^{p}_{\omega}|\mathcal{F}_{s}}+\big\||\varphi_{r}-{\mathbb{E}}^{s}\varphi_{r}|^{\alpha}\big\|_{L^{p}_{\omega}|\mathcal{F}_{s}})\mathrm{d}r.$
-Using (3.7) and the induction hypothesis, we get
-$\|\varphi_{t}-{\mathbb{E}}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}\leqslant N|t-s|^{(H\alpha)\wedge(m\alpha)+1}\quad a.s.$
-We note that $m_{0}=0,m_{i+1}=1+(H\alpha)\wedge(m_{i}\alpha)$ reaches $1+H\alpha$ in finitely many steps, therefore we get (4.1).
-Inequality (4.2) is obtained by a similar argument.
+We only give the proof of (4.2).
 For fixed $s,$ define $s^{\prime}$ to be the smallest grid point which is bigger or equal to $s,$ that is, $s^{\prime}:=\lceil ns\rceil n^{-1}$.
 It is crucial to note that $\varphi_{s^{\prime}}^{n}$ is $\mathcal{F}_{s}$-measurable.
-We proceed by induction as before and assume that (4.2) holds for some $m\geqslant 0.$ If $s\leqslant t<s^{\prime},$ then $\varphi_{t}^{n}$ is $\mathcal{F}_{s}$-measurable.
+Suppose (4.2) holds for some $m\geqslant 0$ in place of $1+\alpha H.$ This is certainly true for $m=0$ thanks to the fact that $b$ is bounded; we proceed now by induction on $m.$ If $s\leqslant t<s^{\prime},$ then $\varphi_{t}^{n}$ is $\mathcal{F}_{s}$-measurable.
 Hence $\varphi^{n}_{t}={\mathbb{E}}^{s}\varphi_{t}^{n}$ and the left-hand side of (4.2) is zero.
-Therefore it remains to consider the case $t\geqslant s^{\prime}.$ In this case, using again (2.2) with $X=\varphi_{t}^{n},Y=\varphi^{n}_{s^{\prime}}+\int_{s^{\prime}}^{t}b({\mathbb{E}}^{s}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s}\varphi_{k_{n}(r)}^{n})\mathrm{d}r,$ we deduce
+Therefore it remains to consider the case $t\geqslant s^{\prime}.$ In this case, using (2.3) with $X=\varphi_{t}^{n},Y=\varphi^{n}_{s^{\prime}}+\int_{s^{\prime}}^{t}b({\mathbb{E}}^{s}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s}\varphi_{k_{n}(r)}^{n})\mathrm{d}r,$ we deduce
 $\|\varphi_{t}^{n}-{\mathbb{E}}^{s}\varphi_{t}^{n}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant 2\left\|\varphi_{t}^{n}-\varphi^{n}_{s^{\prime}}-\int_{s^{\prime}}^{t}b({\mathbb{E}}^{s}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s}\varphi_{k_{n}(r)}^{n})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
 $=2\left\|\int_{s^{\prime}}^{t}(b(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})-b({\mathbb{E}}^{s}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s}\varphi_{k_{n}(r)}^{n})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
-$\leqslant N\left\|\int_{s^{\prime}}^{t}(|B^{H}_{k_{n}(r)}-{\mathbb{E}}^{s}B^{H}_{k_{n}(r)}|^{\alpha}+|\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s}\varphi^{n}_{k_{n}(r)}|^{\alpha})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}.$
-From now on we obtain (4.2) as before. ∎
+$\leqslant C\|b\|_{{\mathcal{C}}^{\alpha}}\left\|\int_{s^{\prime}}^{t}(|B^{H}_{k_{n}(r)}-{\mathbb{E}}^{s}B^{H}_{k_{n}(r)}|^{\alpha}+|\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s}\varphi^{n}_{k_{n}(r)}|^{\alpha})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}.$
+Using (3.7) and the induction hypothesis, we get
+$\|\varphi_{t}-{\mathbb{E}}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}\leqslant C\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{(H\alpha)\wedge(m\alpha)+1}\quad a.s.$
+We note that $m_{0}=0,m_{i+1}=1+(H\alpha)\wedge(m_{i}\alpha)$ reaches $1+H\alpha$ in finitely many steps, therefore we get (4.1). ∎
 Here we introduce some notations commonly used in the proofs of Lemma˜4.2, Lemma˜4.3 and Lemma˜4.4. For $(s,u,t)\in\overline{[0,1]}_{1}^{3}$, we set
@@
 Note by the fact $u\leqslant\frac{2}{3}s+\frac{1}{3}t$ for $(s,u,t)\in\overline{[0,1]}_{1}^{3}$, we have $s_{2}\leqslant s_{3}.$
-Let us start with the estimate for $\mathcal{E}^{b,n,1}$.
+Let us start with the estimate for $\mathcal{E}^{f,n,1}$.
 Lemma 4.2.
-Suppose Assumption˜2.1 holds. Then for any $p\geqslant 1$ we have
-$\left\|\mathcal{E}^{b,n,1}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant N\|\varphi-\varphi^{n}\|_{C_{p}^{\frac{1}{2}}}|t-s|^{\frac{1}{2}+\varepsilon},\quad\forall(s,t)\in[0,1]_{0}^{2}$ (4.4)
-with sufficiently small $\varepsilon>0$ and some constant $N=N(p,d,\alpha,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
+Suppose Assumption˜2.1 holds. Then for any $p\geqslant 1$ and $\gamma\in(\frac{1}{2},1+(\alpha-1)H)$ we have
+$\left\|\mathcal{E}^{f,n,1}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{\gamma}[s,t]}|t-s|^{\gamma+\varepsilon},\quad(s,t)\in[0,1]_{0}^{2},f\in{\mathcal{C}}^{\alpha}$ (4.4)
+with sufficiently small $\varepsilon>0$ and some constant $C=C(p,d,\alpha,\gamma,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
 Proof.
 The idea is to apply Lemma˜3.1. Let $M=1$, $(s,t)\in[0,1]_{1}^{2}$ and
-$A_{s,t}:={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}b(B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})-b(B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})\mathrm{d}r.$
+$A_{s,t}:={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}f(B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})-f(B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})\mathrm{d}r.$
 We are going to verify (3.1) and (3.2). By (3.7), we see
-$A_{s,t}$ $=\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})$
-$\qquad-\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi^{n}_{r})\mathrm{d}r.$ (4.5)
+$A_{s,t}$ $=\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}f({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})$
+$\qquad-\mathcal{P}^{H}_{r-[s-(t-s)]}f({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi^{n}_{r})\mathrm{d}r.$ (4.5)
 Then by (3.9) and (2.2), we get
-$\|A_{s,t}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(r-[s-(t-s)])^{-(1-\alpha)H}\|{\mathbb{E}}^{s-(t-s)}(\varphi_{r}-\varphi_{r}^{n})\|_{L^{p}_{\omega}}\mathrm{d}r$ (4.6)
-$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}|t-s|^{1-(1-\alpha)H}.$
-Then (3.1) holds with $C_{1}=N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}$ by the fact that $1-(1-\alpha)H>\frac{1}{2}$.
+$\|A_{s,t}\|_{L^{p}_{\omega}}$ $\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(r-[s-(t-s)])^{-(1-\alpha)H}\|{\mathbb{E}}^{s-(t-s)}(\varphi_{r}-\varphi_{r}^{n})\|_{L^{p}_{\omega}}\mathrm{d}r$ (4.6)
+$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}[s,t]}|t-s|^{1-(1-\alpha)H}.$
+Then (3.1) holds with $C_{1}=C\|f\|_{{\mathcal{C}}^{\alpha}[s,t]}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}$ by the fact that $1-(1-\alpha)H>\frac{1}{2}$.
 Next we verify (3.2). Let $(s,u,t)\in\overline{[0,1]}_{1}^{3}$. Recall the definition of $s_{i},i=1,\dots,6$ in (4.3). We first can write
 ${\mathbb{E}}^{s-(t-s)}\delta A_{s,u,t}$ (4.7)
-$=$ ${\mathbb{E}}^{s_{1}}{\mathbb{E}}^{s_{3}}\int_{s_{4}}^{s_{5}}b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-b(B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r})+b(B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})\mathrm{d}r$
-$+{\mathbb{E}}^{s_{1}}{\mathbb{E}}^{s_{2}}\int_{s_{5}}^{s_{6}}b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-b(B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})+b(B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})\mathrm{d}r$
-$=$ ${\mathbb{E}}^{s_{1}}\int_{s_{4}}^{s_{5}}\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})$
-$\qquad\qquad-\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r})+\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})\mathrm{d}r$
-$+{\mathbb{E}}^{s_{1}}\int_{s_{5}}^{s_{6}}\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})$
-$\qquad\qquad-\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})+\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})\mathrm{d}r$
+$=$ ${\mathbb{E}}^{s_{1}}{\mathbb{E}}^{s_{3}}\int_{s_{4}}^{s_{5}}f(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-f(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-f(B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r})+f(B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})\mathrm{d}r$
+$+{\mathbb{E}}^{s_{1}}{\mathbb{E}}^{s_{2}}\int_{s_{5}}^{s_{6}}f(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-f(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-f(B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})+f(B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})\mathrm{d}r$
+$=$ ${\mathbb{E}}^{s_{1}}\int_{s_{4}}^{s_{5}}\mathcal{P}^{H}_{r-s_{3}}f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{3}}f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})$
+$\qquad\qquad-\mathcal{P}^{H}_{r-s_{3}}f({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r})+\mathcal{P}^{H}_{r-s_{3}}f({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})\mathrm{d}r$
+$+{\mathbb{E}}^{s_{1}}\int_{s_{5}}^{s_{6}}\mathcal{P}_{r-s_{2}}^{H}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{2}}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})$
+$\qquad\qquad-\mathcal{P}^{H}_{r-s_{2}}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})+\mathcal{P}^{H}_{r-s_{3}}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})\mathrm{d}r$
 $=:$ $I_{1}+I_{2}.$
@@
 we obtain
-$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}|r-s_{3}|^{-(1-\alpha)H}\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}$ (4.8)
+$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}|r-s_{3}|^{-(1-\alpha)H}\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}$ (4.8)
 $\quad\qquad\qquad\quad+|r-s_{3}|^{-(2-\alpha)H}\big\||{\mathbb{E}}^{s_{1}}\varphi_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\cdot{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}\mathrm{d}r.$
 By (4.2),
-${\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|={\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}(\varphi^{n}_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})|\leqslant{\mathbb{E}}^{s_{1}}|\varphi^{n}_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\leqslant N|r-s_{1}|^{1+\alpha H}.$ (4.9)
+${\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|={\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}(\varphi^{n}_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})|\leqslant{\mathbb{E}}^{s_{1}}|\varphi^{n}_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\leqslant C|r-s_{1}|^{1+\alpha H}.$ (4.9)
 Besides from (2.2), we get
@@
 meanwhile (2.4) implies
-$\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}\leqslant N|r-s_{1}|^{\frac{1}{2}}[\varphi-\varphi^{n}]_{C^{\frac{1}{2}}_{p}},$ (4.11)
+$\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}\leqslant C|r-s_{1}|^{\gamma}[\varphi-\varphi^{n}]_{C^{\gamma}_{p}[s,t]},$ (4.11)
 clearly from (2.2),
-$\|{\mathbb{E}}^{s_{1}}(\varphi_{r}-\varphi_{r}^{n})\|_{L^{p}_{\omega}}\leqslant\|\varphi-\varphi^{n}\|_{C_{p}^{0}},$ (4.12)
+$\|{\mathbb{E}}^{s_{1}}(\varphi_{r}-\varphi_{r}^{n})\|_{L^{p}_{\omega}}\leqslant\|\varphi-\varphi^{n}\|_{C_{p}^{0}[s,t]},$ (4.12)
 now plugging (4.9), (4.11) and (4.12) into (4.8), we have
-$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}[\varphi-\varphi^{n}]_{C_{p}^{\frac{1}{2}}}\int_{s_{4}}^{s_{5}}(r-s_{1})^{\frac{1}{2}}(r-s_{3})^{-(1-\alpha)H}\mathrm{d}r$ (4.13)
-$\quad+\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}\int_{s_{4}}^{s_{5}}(r-s_{1})^{1+\alpha H}(r-s_{3})^{-(2-\alpha)H}\mathrm{d}r$
-$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{\frac{3}{2}-(1-\alpha)H}[\varphi-\varphi^{n}]_{C_{p}^{\frac{1}{2}}}$
-$\quad+\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{2+\alpha H-(2-\alpha)H}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}$
-$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{\frac{1}{2}}}(t-s)^{(\frac{3}{2}-(1-\alpha)H)\wedge(2+\alpha H-(2-\alpha)H)}.$
+$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}[\varphi-\varphi^{n}]_{C_{p}^{\gamma}[s,t]}\int_{s_{4}}^{s_{5}}(r-s_{1})^{\gamma}(r-s_{3})^{-(1-\alpha)H}\mathrm{d}r$ (4.13)
+$\quad+\|f\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}[s,t]}\int_{s_{4}}^{s_{5}}(r-s_{1})^{1+\alpha H}(r-s_{3})^{-(2-\alpha)H}\mathrm{d}r$
+$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}(t-s)^{1+\gamma-(1-\alpha)H}[\varphi-\varphi^{n}]_{C_{p}^{\gamma}[s,t]}$
+$\quad+\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{2+\alpha H-(2-\alpha)H}\|\varphi-\varphi^{n}\|_{C_{p}^{0}[s,t]}$
+$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{\gamma}[s,t]}(t-s)^{(1+\gamma-(1-\alpha)H)\wedge(2+\alpha H-(2-\alpha)H)}.$
 The above analysis also implies the same bound on $I_{2}$ observing $I_{1}$ and $I_{2}$ share the same structure.
-Noticing Assumption˜2.1 implies $(\frac{3}{2}-(1-\alpha)H)\wedge(2+\alpha H-(2-\alpha)H)>1$, we conclude (3.2) holds with $C_{2}=N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{\frac{1}{2}}}$.
+Noticing Assumption˜2.1 implies $(1+\gamma-(1-\alpha)H)\wedge(2+\alpha H-(2-\alpha)H)>1$, we conclude (3.2) holds with $C_{2}=C\|f\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{\gamma}[s,t]}$.
 Now we claim that the process $\mathcal{A}$ in (3.3) actually is given by
-${\mathcal{A}}_{t}=\int_{0}^{t}b(B^{H}_{r}+\varphi_{r})-b(B^{H}_{r}+\varphi_{r}^{n})\mathrm{d}r.$ (4.14)
+${\mathcal{A}}_{t}=\int_{0}^{t}f(B^{H}_{r}+\varphi_{r})-f(B^{H}_{r}+\varphi_{r}^{n})\mathrm{d}r.$ (4.14)
 To prove this, it suffices to show (3.4). By (4), we write
-${\mathcal{A}}_{t}-{\mathcal{A}}_{s}-A_{s,t}$ $=\int_{s}^{t}b(B^{H}_{r}+\varphi_{r})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})\mathrm{d}r$ (4.15)
-$\quad-\int_{s}^{t}b(B_{r}^{H}+\varphi_{r}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})\mathrm{d}r$
+${\mathcal{A}}_{t}-{\mathcal{A}}_{s}-A_{s,t}$ $=\int_{s}^{t}f(B^{H}_{r}+\varphi_{r})-\mathcal{P}^{H}_{r-s_{1}}f({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})\mathrm{d}r$ (4.15)
+$\quad-\int_{s}^{t}f(B_{r}^{H}+\varphi_{r}^{n})-\mathcal{P}^{H}_{r-s_{1}}f({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})\mathrm{d}r$
 $=:II_{1}+II_{2}.$
 Again we can see that $II_{2}$ can be treated similarly to $II_{1}$, so we only detail $II_{1}$. We can see
-$II_{1}$ $=\int_{s}^{t}(b-\mathcal{P}^{H}_{r-s_{1}}b)(B_{r}^{H}+\varphi_{r})\mathrm{d}r+\int_{s}^{t}\mathcal{P}^{H}_{r-s_{1}}b(B_{r}^{H}+\varphi_{r})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})\mathrm{d}r.$ (4.16)
+$II_{1}$ $=\int_{s}^{t}(f-\mathcal{P}^{H}_{r-s_{1}}f)(B_{r}^{H}+\varphi_{r})\mathrm{d}r+\int_{s}^{t}\mathcal{P}^{H}_{r-s_{1}}f(B_{r}^{H}+\varphi_{r})-\mathcal{P}^{H}_{r-s_{1}}f({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})\mathrm{d}r.$ (4.16)
 Using (3.10) with $\delta=\frac{\alpha}{2}$, $\beta=0$ and (3.9) with $\beta=\alpha$, we get
-$\|II_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\int_{s}^{t}\|\mathcal{P}^{H}_{r-s_{1}}b-b\|_{{\mathcal{C}}^{0}}+\|\mathcal{P}^{H}_{r-s_{1}}b\|_{{\mathcal{C}}^{\alpha}}(\|B_{r}^{H}-{\mathbb{E}}^{s_{1}}B^{H}_{r}\|^{\alpha}_{L^{\alpha p}_{\omega}}+\|\varphi_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}\|^{\alpha}_{L_{\omega}^{\alpha p}})\mathrm{d}r$
-$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(r-s_{1})^{\alpha H}+(|r-s_{1}|^{\alpha H}+|r-s_{1}|^{\alpha(1+\alpha H)})\mathrm{d}r$
-$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{(1+\alpha H)\wedge(1+\alpha(1+\alpha H))},$ (4.17)
+$\|II_{1}\|_{L^{p}_{\omega}}$ $\leqslant C\int_{s}^{t}\|\mathcal{P}^{H}_{r-s_{1}}f-f\|_{{\mathcal{C}}^{0}}+\|\mathcal{P}^{H}_{r-s_{1}}f\|_{{\mathcal{C}}^{\alpha}}(\|B_{r}^{H}-{\mathbb{E}}^{s_{1}}B^{H}_{r}\|^{\alpha}_{L^{\alpha p}_{\omega}}+\|\varphi_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}\|^{\alpha}_{L_{\omega}^{\alpha p}})\mathrm{d}r$
+$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(r-s_{1})^{\alpha H}+(|r-s_{1}|^{\alpha H}+|r-s_{1}|^{\alpha(1+\alpha H)})\mathrm{d}r$
+$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}(t-s)^{(1+\alpha H)\wedge(1+\alpha(1+\alpha H))},$ (4.17)
 where in the second inequality we used (3.6) and (4.1). The same bound on $II_{2}$.
 Therefore, (3.4) holds since $(1+\alpha H)\wedge(1+\alpha(1+\alpha H))>1$. Then the uniqueness from Lemma˜3.1 verifies the claim.
-Finally, by Lemma˜3.1, the proof is complete. ∎
-Let us move to estimate $\mathcal{E}^{b,n,2}$ term.
+Finally, by Lemma˜3.1, the proof completes. ∎
+Let us move to estimate $\mathcal{E}^{f,n,2}$ term.
 Lemma 4.3.
-Suppose Assumption˜2.1 holds. Then we have for any $p\geqslant 1$ and $(s,t)\in[0,1]_{0}^{2}$
-$\left\|\mathcal{E}^{b,n,2}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}|t-s|^{\frac{1}{2}+\varepsilon}$ (4.18)
-where $\epsilon>0$ is sufficiently small and $N=N(p,d,\alpha,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
+Suppose Assumption˜2.1 holds. Then we have for any $p\geqslant 1$, $\gamma\in(\frac{1}{2},1+(\alpha-1)H)$ and $(s,t)\in[0,1]_{0}^{2}$
+$\left\|\mathcal{E}^{f,n,2}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\gamma+\varepsilon},\quad(s,t)\in[0,1]_{0}^{2},f\in{\mathcal{C}}^{\alpha}$ (4.18)
+where $\epsilon>0$ is sufficiently small and $C=C(p,d,\alpha,\gamma,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
 Proof.
 Again the idea is to apply Lemma˜3.1. Let $M=1$, $(s,t)\in[0,1]_{1}^{2}$ and
-$A_{s,t}:={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}b(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})-b(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$ (4.19)
-Similarly to (4) and (4.6), we have
-$\|A_{s,t}\|_{L^{p}_{\omega}}$ $=\|\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})$
-$\qquad-\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi^{n}_{k_{n}(r)})\mathrm{d}r\|_{L^{p}_{\omega}}$ (4.20)
-$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\sup_{r\in[s,t]}\|\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}\|_{L_{\omega}^{p}}|t-s|^{1-(1-\alpha)H}$
-$\leqslant Nn^{-1}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1-(1-\alpha)H},$ (4.21)
-where in the second inequality we used $\|\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}\|_{L_{\omega}^{p}}\leqslant\|b\|_{{\mathcal{C}}^{0}}n^{-1}$.
-Then, with taking $C_{1}=N\|b\|_{{\mathcal{C}}^{\alpha}}n^{-1}$, (3.1) holds by the fact that $1-(1-\alpha)H>\frac{1}{2}.$
-Next we verify (3.2). Let $(s,u,t)\in\overline{[0,1]}_{1}^{3}$. Recall the definition of $s_{i},i=1,\dots,6$ in (4.3). Similarly to (4.7), we can write
-${\mathbb{E}}^{s_{1}}\delta A_{s,u,t}=$ ${\mathbb{E}}^{s_{1}}\int_{s_{4}}^{s_{5}}\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
-$\qquad\qquad-\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})+\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
-$+{\mathbb{E}}^{s_{1}}\int_{s_{5}}^{s_{6}}\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
-$\qquad\qquad-\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})+\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
-$=$ $:I_{1}+I_{2}.$
-The two terms are treated in the exactly same manner, so we only detail $I_{1}$. Similarly to (4.8), we get
-$\|I_{1}\|_{L_{\omega}^{p}}\leqslant$ $N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(1-\alpha)H}\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\big\|_{L_{\omega}^{p}}$
-$\qquad+(r-s_{3})^{-(2-\alpha)H}\big\||{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\cdot{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\big\|_{L_{\omega}^{p}}\mathrm{d}r.$ (4.22)
-Similar to (4.10), we have
-$\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\|_{L^{p}_{\omega}}\leqslant\|{\mathbb{E}}^{s_{1}}|(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})|\|_{L^{p}_{\omega}}.$ (4.23)
-We note that
-$\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}$ $=\int_{k_{n}(r)}^{r}b(B^{H}_{k_{n}(t)}+\varphi_{k_{n}(t)}^{n})\mathrm{d}t=(r-k_{n}(r))b(B_{k_{n}(r)}^{H}+\varphi^{n}_{k_{n}(r)})\in\mathcal{F}_{k_{n}(r)}.$
-When $s_{1}\in[k_{n}(r),r],$ we have
-$\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})=0;$
-[…]
-${\mathbb{E}}^{s_{1}}|(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})|$
-$=$ $(r-k_{n}(r)){\mathbb{E}}^{s_{1}}|b(B_{k_{n}(r)}^{H}+\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}b(B_{k_{n}(r)}^{H}+\varphi_{k_{n}(r)}^{n})|$
-$\leqslant$ $2(r-k_{n}(r)){\mathbb{E}}^{s_{1}}|b(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})-b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+\varphi_{s_{1}}^{n})|$
-$\leqslant$ $N(r-k_{n}(r))\|b\|_{{\mathcal{C}}^{\alpha}}{\mathbb{E}}^{s_{1}}(|B_{k_{n}(r)}^{H}-{\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}|^{\alpha}+|\varphi_{k_{n}(r)}^{n}-\varphi_{s_{1}}^{n}|^{\alpha}).$
-Moreover, using (3.5) and (4.2), we have
-${\mathbb{E}}^{s_{1}}|(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})|$
-$\leqslant$ $N(r-k_{n}(r))\|b\|_{{\mathcal{C}}^{\alpha}}(|k_{n}(r)-s_{1}|^{\alpha H}+|k_{n}(r)-s_{1}|^{\alpha})$
-$\leqslant$ $\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|r-s_{1}|^{\alpha}$
-where we used the fact $H>1$ in the second inequality. Plugging it into (4.23), we get
-$\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|r-s_{1}|^{\alpha}.$ (4.24)
-Meanwhile (4.9) and (4.2) yield
-${\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\leqslant N|k_{n}(r)-s_{1}|^{1+\alpha H}\leqslant N|r-s_{1}|^{1+\alpha H},$ (4.25)
-$\|{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})\|_{L_{\omega}^{p}}\leqslant N\|\varphi_{\cdot}^{n}-\varphi_{k_{n}(\cdot)}^{n}\|_{C_{p}^{0}}\leqslant N\frac{\|b\|_{C^{0}}}{n}.$ (4.26)
-Applying (4.24), (4.25) and (4.26) into (4.22) gives us
-$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\frac{\|b\|_{{\mathcal{C}}^{\alpha}}}{n}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(1-\alpha)H}(r-s_{1})^{\alpha}\mathrm{d}r+(r-s_{3})^{-(2-\alpha)H}(r-s_{1})^{1+\alpha H}\mathrm{d}r$
-$\leqslant N\frac{\|b\|_{{\mathcal{C}}^{\alpha}}}{n}(t-s)^{(1+\alpha-(1-\alpha)H)\wedge(2+(2\alpha-2)H)}.$ (4.27)
-With the same bound on $I_{2}$ and Assumption˜2.1 implying $(1+\alpha-(1-\alpha)H)\wedge(2+(2\alpha-2)H)>1$, (3.2) holds with $C_{2}=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}$.
-Lastly we claim that the process $\mathcal{A}$ in (3.3) is indeed given by
-$\mathcal{A}_{t}=\int_{0}^{t}b(B^{H}_{r}+\varphi_{r}^{n})-b(B^{H}_{r}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$ (4.28)
-To prove this claim, it suffices to show (3.4). Similarly to (4), we have
-$\mathcal{A}t-\mathcal{A}_{s}-A_{s,t}$ $=\int_{s}^{t}b(B^{H}_{r}+\varphi_{r}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi^{n}_{r})\mathrm{d}r$
-$\quad-\int_{s}^{t}b(B_{r}^{H}+\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$
-Then the estimates can be done similarly to that in Lemma˜4.2.
-Therefore, the claim holds by the uniqueness from Lemma˜3.1. Finally, by Lemma˜3.1, we can conclude (4.18). ∎
+$A_{s,t}:={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}f(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})-f(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$ (4.19)
+We may notice that the analysis of showing (4.6) and estimating (4.7) also work for (4.19).
+Therefore we omit the detailed proof here for showing the following estimates and present in Appendix˜A; in the end we get
+$\|A_{s,t}\|_{L^{p}_{\omega}}$ $\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\sup_{r\in[s,t]}\|\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}\|_{L_{\omega}^{p}}|t-s|^{1-(1-\alpha)H}$
+$\leqslant Cn^{-1}\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1-(1-\alpha)H},$ (4.20)
+$\|{\mathbb{E}}^{s_{1}}\delta A_{s,u,t}\|_{L^{p}_{\omega}}$ $\leqslant C\frac{\|f\|_{{\mathcal{C}}^{\alpha}}}{n}(t-s)^{(1+\alpha-(1-\alpha)H)\wedge(2+(2\alpha-2)H)},$ (4.21)
+recall the definition of $s_{i},i=1,\dots,6$ in (4.3).
+Then, with taking $C_{1}=C\|f\|_{{\mathcal{C}}^{\alpha}}n^{-1}$, (3.1) holds by the fact that $1-(1-\alpha)H>\frac{1}{2}.$ Assumption˜2.1 implying $(1+\alpha-(1-\alpha)H)\wedge(2+(2\alpha-2)H)>1$, (3.2) holds with $C_{2}=\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}$
+Similarly to Lemma˜4.2, we could verify that the process $\mathcal{A}$ in (3.3) is indeed given by
+$\mathcal{A}_{t}=\int_{0}^{t}f(B^{H}_{r}+\varphi_{r}^{n})-f(B^{H}_{r}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$ (4.22)
+Therefore, the claim (4.18) holds by the uniqueness from Lemma˜3.1. ∎
 […]
 Lemma 4.4.
-Suppose Assumption˜2.1 holds. Then for any $p\geqslant 1$ and $(s,t)\in[0,1]_{0}^{2}$, we have
-$\left\|\mathcal{E}^{b,n,3}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}|t-s|^{\frac{1}{2}+\varepsilon}$ (4.29)
-with sufficiently small $\varepsilon>0$ and some constant $N=N(p,d,\alpha,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
+Suppose Assumption˜2.1 holds. Then for any $p\geqslant 1$, $\gamma\in(\frac{1}{2},1+(\alpha-1)H)$ and $(s,t)\in[0,1]_{0}^{2}$, we have
+$\left\|\mathcal{E}^{f,n,3}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\gamma+\varepsilon},\quad(s,t)\in[0,1]_{0}^{2},f\in{\mathcal{C}}^{\alpha}$ (4.23)
+with sufficiently small $\varepsilon>0$ and some constant $C=C(p,d,\alpha,\gamma,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
 Proof.
 In order to apply Lemma˜3.1, this time we set
-$A_{s,t}={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}b(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})-b(B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$
+$A_{s,t}={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}f(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})-f(B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$
 When $|t-s|\leqslant\frac{10}{n}$, by (3.6), we have for any $\epsilon>0$
-$\|A_{s,t}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}\big\||B^{H}_{r}-B^{H}_{k_{n}(r)}|^{\alpha}\big\|_{L^{p}_{\omega}}\mathrm{d}r\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|\cdot\frac{1}{n^{\alpha}}$
-$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\frac{1}{2}+\varepsilon}\frac{1}{n^{\alpha+\frac{1}{2}-\varepsilon}}.$
-For $H>1$ and $\alpha>1-\frac{1}{2H}$ we can take such sufficiently small $\varepsilon>0$ so that $\epsilon\in(0,\alpha+\frac{1}{2H}-1)$, which implies
-$\|A_{s,t}\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\frac{1}{2}+\varepsilon}.$ (4.30)
+$\|A_{s,t}\|_{L^{p}_{\omega}}$ $\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}\big\||B^{H}_{r}-B^{H}_{k_{n}(r)}|^{\alpha}\big\|_{L^{p}_{\omega}}\mathrm{d}r\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|\cdot\frac{1}{n^{\alpha}}$
+$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\gamma+\varepsilon}\frac{1}{n^{\alpha+1-\gamma-\varepsilon}}.$
+Since $\alpha>\gamma$ we can take such sufficiently small $\varepsilon>0$ so that $\epsilon\in(0,\alpha-\gamma)$, which implies
+$\|A_{s,t}\|_{L^{p}_{\omega}}\leqslant\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\frac{1}{2}+\varepsilon}.$ (4.24)
 When $|t-s|>\frac{10}{n}$, we have
-$A_{s,t}$ $=\int_{s_{4}}^{s_{6}}\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{k_{n}(r)-s_{1}}b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
-$=\int_{s_{4}}^{s_{6}}\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
-$\qquad+\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{k_{n}(r)-s_{1}}b({\mathbb{E}}^{s_{1}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
+$A_{s,t}$ $=\int_{s_{4}}^{s_{6}}\mathcal{P}^{H}_{r-s_{1}}f({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{k_{n}(r)-s_{1}}f({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
+$=\int_{s_{4}}^{s_{6}}\mathcal{P}^{H}_{r-s_{1}}f({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{r-s_{1}}f({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
+$\qquad+\mathcal{P}^{H}_{r-s_{1}}f({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{k_{n}(r)-s_{1}}f({\mathbb{E}}^{s_{1}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
 $=:IV_{1}+IV_{2}.$
 For $IV_{1}$, by (3.9), (2.2) and (3.6), we have
-$\|IV_{1}\|_{L^{p}_{\omega}}$ $\leqslant\int_{s_{4}}^{s_{6}}\|\mathcal{P}^{H}_{r-s_{1}}b\|_{{\mathcal{C}}^{1}}\|{\mathbb{E}}^{s_{1}}(B_{r}^{H}-B_{k_{n}(r)}^{H})\|_{L^{p}_{\omega}}\mathrm{d}r$
-$\leqslant\frac{N\|b\|_{{\mathcal{C}}^{\alpha}}}{n}\int_{s_{4}}^{s_{6}}(r-s_{1})^{-(1-\alpha)H}\mathrm{d}r$
-$\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1-(1-\alpha)H}.$ (4.31)
+$\|IV_{1}\|_{L^{p}_{\omega}}$ $\leqslant\int_{s_{4}}^{s_{6}}\|\mathcal{P}^{H}_{r-s_{1}}f\|_{{\mathcal{C}}^{1}}\|{\mathbb{E}}^{s_{1}}(B_{r}^{H}-B_{k_{n}(r)}^{H})\|_{L^{p}_{\omega}}\mathrm{d}r\leqslant\frac{C\|f\|_{{\mathcal{C}}^{\alpha}}}{n}\int_{s_{4}}^{s_{6}}(r-s_{1})^{-(1-\alpha)H}\mathrm{d}r$
+$\leqslant\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1-(1-\alpha)H}.$ (4.25)
 For $IV_{2}$, (3.10) with $\delta=1$ gives us
-$\|\mathcal{P}^{H}_{t}b-\mathcal{P}_{s}^{H}b\|_{{\mathcal{C}}^{0}}\leqslant Ns^{-(2-\alpha)H}|t^{2H}-s^{2H}|\|b\|_{{\mathcal{C}}^{\alpha}},$
+$\|\mathcal{P}^{H}_{t}f-\mathcal{P}_{s}^{H}f\|_{{\mathcal{C}}^{0}}\leqslant Cs^{-(2-\alpha)H}|t^{2H}-s^{2H}|\|f\|_{{\mathcal{C}}^{\alpha}},$
 it implies
-$|IV_{2}|$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{6}}(k_{n}(r)-s_{1})^{-(2-\alpha)H}\left((r-s_{1})^{2H}-(k_{n}(r)-s_{1})^{2H}\right)\mathrm{d}r.$
+$|IV_{2}|$ $\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{6}}(k_{n}(r)-s_{1})^{-(2-\alpha)H}\left((r-s_{1})^{2H}-(k_{n}(r)-s_{1})^{2H}\right)\mathrm{d}r.$
 Moreover, by $k_{n}(r)-s_{1}\asymp t-s,$ and
-$\big||r-s_{1}|^{2H}-|k_{n}(r)-s_{1}|^{2H}\big|\leqslant N|r-k_{n}(r)||r-s_{1}|^{2H-1}\leqslant\frac{N}{n}|t-s|^{2H-1},$
+$\big||r-s_{1}|^{2H}-|k_{n}(r)-s_{1}|^{2H}\big|\leqslant C|r-k_{n}(r)||r-s_{1}|^{2H-1}\leqslant\frac{C}{n}|t-s|^{2H-1},$
 we have
-$|IV_{2}|\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{6}}(t-s)^{2H-1-(2-\alpha)H}\mathrm{d}r=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{\alpha H}.$ (4.32)
-Then (4.30) togehter with (4.31) and (4.32) verifies (3.1) of Lemma˜3.1 with taking $C_{1}=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}$, since $((1-(1-\alpha)H)\wedge(\alpha H\big))>\frac{1}{2}$.
+$|IV_{2}|\leqslant\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{6}}(t-s)^{2H-1-(2-\alpha)H}\mathrm{d}r=\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}(t-s)^{\alpha H}.$ (4.26)
+Then (4.24) togehter with (4.25) and (4.26) verifies (3.1) of Lemma˜3.1 with taking $C_{1}=\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}$, since $((1-(1-\alpha)H)\wedge(\alpha H\big))>\gamma>\frac{1}{2}$.
 Next we verify (3.2). Let $(s,u,t)\in\overline{[0,1]}_{1}^{3}$. Recall the definition of $s_{i},i=1,\dots,6$ in (4.3). Similarly to (4.7), we have
 […]
-$\qquad-\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n})+\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
+$\qquad-\mathcal{P}_{r-s_{3}}^{H}f({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n})+\mathcal{P}^{H}_{r-s_{3}}f({\mathbb{E}}^{s_{3}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
 […]
-$\qquad-\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B^{H}_{r}+{\mathbb{E}}^{s_{2}}\varphi_{k_{n}(r)}^{n})+\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{2}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
+$\qquad-\mathcal{P}^{H}_{r-s_{2}}f({\mathbb{E}}^{s_{2}}B^{H}_{r}+{\mathbb{E}}^{s_{2}}\varphi_{k_{n}(r)}^{n})+\mathcal{P}^{H}_{r-s_{2}}f({\mathbb{E}}^{s_{2}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{2}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
 $=:V_{1}+V_{2}.$
@@
 $\|V_{1}\|_{L^{p}_{\omega}}$
-$\leqslant N\int_{s_{4}}^{s_{5}}\|b\|_{{\mathcal{C}}^{\alpha}}(r-s_{3})^{-H(2-\alpha)}\|{\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{3}}(B_{r}^{H}-B_{k_{n}(r)}^{H})|\cdot|{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)}|]\|_{L^{p}_{\omega}}\mathrm{d}r.$ (4.33)
+$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-H(2-\alpha)}\|{\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{3}}(B_{r}^{H}-B_{k_{n}(r)}^{H})|\cdot|{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)}|]\|_{L^{p}_{\omega}}\mathrm{d}r.$ (4.27)
 By Cauchy-Schwarz inequality,
@@
 $\leqslant({\mathbb{E}}^{s_{1}}(B^{H}_{r}-B^{H}_{k_{n}(r)})^{2})^{\frac{1}{2}}\cdot({\mathbb{E}}^{s_{1}}(\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)})^{2})^{\frac{1}{2}}$
-$\leqslant N({\mathbb{E}}^{s_{1}}(B^{H}_{r}-B^{H}_{k_{n}(r)})^{2})^{\frac{1}{2}}\cdot(k_{n}(r)-s_{1})^{1+\alpha H}.$
+$\leqslant C({\mathbb{E}}^{s_{1}}(B^{H}_{r}-B^{H}_{k_{n}(r)})^{2})^{\frac{1}{2}}\cdot(k_{n}(r)-s_{1})^{1+\alpha H}.$
 For $p\geqslant 2$, (2.2) and (3.6) imply
 $\|{\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{3}}(B_{r}^{H}-B^{H}_{k_{n}(r)})||{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)}|]\|_{L^{p}_{\omega}}$ (4.28)
-$\leqslant N(r-s_{1})^{1+\alpha H}\left\|\|B_{r}^{H}-B^{H}_{k_{n}(r)}\|_{L_{\omega}^{2}|\mathcal{F}_{s_{1}}}\right\|_{L^{p}_{\omega}}$
-$=N(r-s_{1})^{1+\alpha H}\|{\mathbb{E}}^{s_{1}}|B_{r}^{H}-B_{k_{n}(r)}^{H}|^{2}\|^{\frac{1}{2}}_{L_{\omega}^{\frac{p}{2}}}$
-$\leqslant N(r-s_{1})^{1+\alpha H}\|B_{r}^{H}-B^{H}_{k_{n}(r)}\|_{L^{p}_{\omega}}$
-$\leqslant\frac{N}{n}(r-s_{1})^{1+\alpha H}.$
+$\leqslant C(r-s_{1})^{1+\alpha H}\left\|\|B_{r}^{H}-B^{H}_{k_{n}(r)}\|_{L_{\omega}^{2}|\mathcal{F}_{s_{1}}}\right\|_{L^{p}_{\omega}}$
+$=C(r-s_{1})^{1+\alpha H}\|{\mathbb{E}}^{s_{1}}|B_{r}^{H}-B_{k_{n}(r)}^{H}|^{2}\|^{\frac{1}{2}}_{L_{\omega}^{\frac{p}{2}}}$
+$\leqslant C(r-s_{1})^{1+\alpha H}\|B_{r}^{H}-B^{H}_{k_{n}(r)}\|_{L^{p}_{\omega}}$
+$\leqslant\frac{C}{n}(r-s_{1})^{1+\alpha H}.$
 Therefore, plugging (4.28) into (4), we have
-$\|V_{1}\|_{L^{p}_{\omega}}$ $\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(2-\alpha)H}(r-s_{1})^{1+\alpha H}\mathrm{d}r\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{2+(2\alpha-2)H}.$
+$\|V_{1}\|_{L^{p}_{\omega}}$ $\leqslant\frac{C}{n}\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(2-\alpha)H}(r-s_{1})^{1+\alpha H}\mathrm{d}r\leqslant\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{2+(2\alpha-2)H}.$
 The same bound holds on $V_{2}$.
-Since Assumption˜2.1 implies $2+(2\alpha-2)H>1$, we can obtain that (3.2) holds with $C_{2}=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}$.
-In order to show that the process $\mathcal{A}$ in (3.3) is given by
-$\mathcal{A}_{t}=\int_{0}^{t}b(B^{H}_{r}+\varphi^{n}_{k_{n}(r)})-b(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r,$
-it suffices to show (3.4). For doing so, we first write
-$\mathcal{A}_{t}-\mathcal{A}_{s}-A_{s,t}$ $=\int_{s}^{t}b(B_{r}^{H}+\varphi^{n}_{k_{n}(r)})-\mathcal{P}_{r-s_{1}}^{H}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
-$\quad-\int_{s}^{t}b(B^{H}_{k_{n}(r)}+\varphi^{n}_{k_{n}(r)})-\mathcal{P}^{H}_{(k_{n}(r)-s_{1})\vee 0}b({\mathbb{E}}^{s_{1}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r.$
-Then similar to what we have done in Lemma˜4.2 (that is, (4), (4.16) and (4.17)), we can conclude the claim by the uniqueness from Lemma˜3.1.
+Since Assumption˜2.1 implies $2+(2\alpha-2)H>1$, we can obtain that (3.2) holds with $C_{2}=\frac{C}{n}\|b\|_{{\mathcal{C}}^{\alpha}}$.
+Similarly to Lemma˜4.2, we could verify the process $\mathcal{A}$ in (3.3) is given by
+$\mathcal{A}_{t}=\int_{0}^{t}f(B^{H}_{r}+\varphi^{n}_{k_{n}(r)})-f(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$
 In the end all of the conditions from Lemma˜3.1 are verified, which proves the desired result. ∎
@@
 By (4.4), (4.18) and (4.23), we see that
-$\|(X-X^{n})_{t}-(X-X^{n})_{s}\|_{L^{p}_{\omega}}$ $\leqslant N(\|\varphi-\varphi^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon}$
-$=N(\|X-X^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon},$
+$\|(X-X^{n})_{t}-(X-X^{n})_{s}\|_{L^{p}_{\omega}}$ $\leqslant C(\|\varphi-\varphi^{n}\|_{C^{\gamma}_{p}[s,t]}+n^{-1})|t-s|^{\gamma+\varepsilon}$
+$=C(\|X-X^{n}\|_{C^{\gamma}_{p}[s,t]}+n^{-1})|t-s|^{\gamma+\varepsilon},$
 which implies
-$[X-X^{n}]_{C_{p}^{\frac{1}{2}}[S,T]}\leqslant N(\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[S,T]}+n^{-1})(T-S)^{\varepsilon}.$
+$[X-X^{n}]_{C_{p}^{\gamma}[S,T]}\leqslant C(\|X-X^{n}\|_{C_{p}^{\gamma}[S,T]}+n^{-1})(T-S)^{\varepsilon}.$
 Therefore, we have
-$\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[S,T]}$ $\leqslant|(X-X^{n})_{S}|+2[X-X^{n}]_{C_{p}^{\frac{1}{2}}[S,T]}$
-$\leqslant|(X-X^{n})_{S}|+N(\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[S,T]}+n^{-1})(T-S)^{\varepsilon}.$
+$\|X-X^{n}\|_{C_{p}^{\gamma}[S,T]}$ $\leqslant|(X-X^{n})_{S}|+2[X-X^{n}]_{C_{p}^{\gamma}[S,T]}$
+$\leqslant|(X-X^{n})_{S}|+C(\|X-X^{n}\|_{C_{p}^{\gamma}[S,T]}+n^{-1})(T-S)^{\varepsilon}.$
 Fix $T-S=\Delta$ small enough and we obtain
-$\|X-X^{n}\|_{C^{\frac{1}{2}}_{p}[S,T]}\leqslant N(|(X-X^{n})_{S}|+n^{-1}).$
+$\|X-X^{n}\|_{C^{\gamma}_{p}[S,T]}\leqslant C(|(X-X^{n})_{S}|+n^{-1}).$
 Dividing $[0,1]$ into $0,\Delta],[\Delta,2\Delta],\ldots$, yields that
-$\|X-X^{n}\|_{C^{\frac{1}{2}}_{p}[0,1]}\leqslant N(|x_{0}-x_{0}^{n}|+n^{-1}).$
+$\|X-X^{n}\|_{C^{\gamma}_{p}[0,1]}\leqslant C(|x_{0}-x_{0}^{n}|+n^{-1}).$
 5. Optimality
-In the following we present a result concerning the optimal convergence rate of EM (1.9) for rather smooth $b\in{\mathcal{C}}^{1}$.
-Theorem 5.1.
-Suppose $b\in{\mathcal{C}}^{1}$ and $x_{0}^{n}=x_{0}$. Then for any $t\in[0,1]$, a.s.
-$\lim_{n\rightarrow\infty}n(X_{t}-X_{t}^{n})=:c(t)$ (5.1)
-exists and $c(t)$ satisfies
-$c^{\prime}(t)=\nabla b(X_{t})c(t)+\frac{1}{2}\nabla b(X_{t})(b(X_{t})+(B_{t}^{H})^{\prime}),\quad c(0)=0.$
+In this section, we discuss the optimality of the convergence rate $n^{-1}$ obtained in the previous part.
+By showing that $n(X-X^{n})$ converges to a non-zero limit, we verify that the rate $n^{-1}$ is optimal for the scheme (1.9).
+In the following first we present a regularization lemma concerning the solution to (1.3).
+Lemma 5.1.
+Let $(X_{t})_{t\in[0,1]}$ be the solution to (1.3). Suppose Assumption˜2.1 holds. Then for every $p>1$, for any $f\in{\mathcal{C}}^{1}$, we have
+$\Big\|\int_{0}^{\cdot}\nabla f(X_{t})\mathrm{d}t\Big\|_{C_{p}^{1+(\alpha-1)H}[0,1]}\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}},$ (5.1)
+where $C$ depends only on $d,p,H,\alpha,\|b\|_{{\mathcal{C}}^{\alpha}}$.
 Proof.
+In order to show (5.1), we apply Lemma˜3.1. Let $M=1$, $(s,t)\in[0,1]_{1}^{2}$ and
+$A_{s,t}:={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}\nabla f(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})\mathrm{d}r\text{ where }\varphi:=X-B^{H}.$
+We are going to verify (3.1) and (3.2). By (3.7), we see
+$A_{s,t}=\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}(\nabla f)({\mathbb{E}}^{s-(t-s)}B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})\mathrm{d}r.$
+Then by (3.9), we have
+$|A_{s,t}|\leqslant\int_{s}^{t}\|\mathcal{P}^{H}_{r-[s-(t-s)]}(\nabla f)\|_{C^{0}}\mathrm{d}r\leqslant C\int_{s}^{t}\left[r-[s-(t-s)]\right]^{H(\alpha-1)}\|f\|_{C^{\alpha}}\mathrm{d}r.$
+Therefore, we have $\|A_{s,t}\|_{L^{p}_{\omega}}\leqslant C\|f\|_{C^{\alpha}}(t-s)^{1-(1-\alpha)H}.$ Then (3.1) holds with $C_{1}=C\|f\|_{C^{\alpha}}$ by the fact that $1-(1-\alpha)H>\frac{1}{2}$.
+Next we verify (3.2). Let $(s,u,t)\in\overline{[0,1]}_{1}^{3}$. Recall the definition of $s_{i},i=1,\dots,6$ in (4.3). We first can write
+${\mathbb{E}}^{s_{1}}\delta A_{s,u,t}$ $={\mathbb{E}}^{s_{1}}\bigg[\int_{s_{4}}^{s_{5}}{\mathbb{E}}^{s_{1}}\nabla f(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-{\mathbb{E}}^{s_{3}}\nabla f(B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{r})\mathrm{d}r\bigg]$
+$\quad+{\mathbb{E}}^{s_{1}}\bigg[\int_{s_{5}}^{s_{6}}{\mathbb{E}}^{s_{1}}\nabla f(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-{\mathbb{E}}^{s_{2}}\nabla f(B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})\mathrm{d}r\bigg]$
+$=\int_{s_{4}}^{s_{5}}{\mathbb{E}}^{s_{1}}\bigg[\mathcal{P}^{H}_{r-s_{3}}\nabla f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{3}}\nabla f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{r})\bigg]\mathrm{d}r$
+$\quad+\int_{s_{5}}^{s_{6}}{\mathbb{E}}^{s_{1}}\bigg[\mathcal{P}^{H}_{r-s_{2}}\nabla f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{2}}\nabla f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})\bigg]\mathrm{d}r.$
 By (3.9) and (4.1), we have
-$n(X_{t}-X_{t}^{n})$ $=\int_{0}^{t}n(b\left(X_{r}\right)-b\left(X_{r}^{n}\right))+n(b\left(X_{r}^{n}\right)-b(X_{k_{n}(r)}^{n}))\mathrm{d}r$
-$=\int_{0}^{t}\int_{0}^{1}\nabla b(X_{r}^{n}+\theta(X_{r}-X_{r}^{n}))n(X_{r}-X_{r}^{n})$
-$\qquad\quad+\nabla b(X_{k_{n}(r)}^{n}+\theta(X_{r}^{n}-X_{k_{n}(r)}^{n}))n(X_{r}^{n}-X_{k_{n}(r)}^{n})\mathrm{d}\theta\mathrm{d}r.$
-By fundamental theory of calculus we can write
-$n(X_{t}-X_{t}^{n})$ $=\int_{0}^{t}\exp\left(\int_{s}^{t}\int_{0}^{1}\nabla b(X_{r}^{n}+\theta(X_{r}-X_{r}^{n}))d\theta dr\right)$
-$\qquad\int_{0}^{1}\nabla b(X_{k_{n}(s)}^{n}+\theta(X_{s}^{n}-X_{k_{n}(s)}^{n}))n(X_{s}^{n}-X_{k_{n}(s)}^{n})\mathrm{d}\theta\mathrm{d}s=:c^{n}(t).$
-Following from (1.9) and the fact a.s.
-$B^{H}\in{\mathcal{C}}^{H-\varepsilon}([0,1])$ for sufficiently small $\varepsilon>0$ by Kolmogorov continuity criteria, which means there exists a small enough $\epsilon^{\prime}>0$ so that a.s.
-$B_{H}\in{\mathcal{C}}^{1+\epsilon^{\prime}}$, we have a.s.
+$\|{\mathbb{E}}^{s_{1}}\delta A_{s,u,t}\|_{L^{p}_{\omega}}$ $\leqslant C\int_{s}^{t}(r-s_{3})^{H(\alpha-2)}\|f\|_{{\mathcal{C}}^{\alpha}}\|{\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{1}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}|]\|_{L^{p}_{\omega}}\mathrm{d}r$
+$\quad+\int_{u}^{t}(r-s_{2})^{H(\alpha-2)}\|f\|_{{\mathcal{C}}^{\alpha}}\|{\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{1}}\varphi_{r}-{\mathbb{E}}^{s_{2}}\varphi_{r}|]\|_{L^{p}_{\omega}}\mathrm{d}r$
+$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{2+2\alpha H-2H}.$
+Noticing Assumption˜2.1 implies $2+2\alpha H-2H>1$, we conclude (3.2) holds with $C_{2}=C\|f\|_{{\mathcal{C}}^{\alpha}}$.
+Moreover, since $f\in C^{1}$, it is obviously to see that the process $\mathcal{A}$ in (3.3) actually is given by
+${\mathcal{A}}_{t}=\int_{0}^{t}\nabla f(X_{r})\mathrm{d}r.$
+Therefore, by Lemma˜3.1, we have
+$\big\|\int_{s}^{t}\nabla f(X_{r})\mathrm{d}r\big\|_{L^{p}}\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{H(\alpha-1)+1},$
+which implies that (5.1) holds. ∎
+Due to (5.1), we can already state a result which is analogous to [7, Proposition 2.8].
+Proposition 5.2.
+Let $(X_{t})_{t\in[0,1]}$ be the solutions to (1.3).
+Suppose Assumption˜2.1 holds.
+For every $f\in{\mathcal{C}}^{\alpha}({\mathbb{R}}^{d})$ there exists a $({\mathcal{F}}_{t})_{t\in[0,1]}$-adapted process $({\mathcal{H}}f)^{X}_{t\in[0,1]}\in C_{p}^{1+(\alpha-1)H-}[0,1]$ for every $p\geqslant 2$ such that for $g\in{\mathcal{C}}^{1}$ with probability one,
+$({\mathcal{H}}g)^{X}_{t}=\int_{0}^{t}\nabla g(X_{s})\mathrm{d}s,\quad t\in[0,1].$
+Proof.
+In fact, we notice that for any $f\in{\mathcal{C}}^{\alpha},$ there exists sequence $(f_{n})_{n\in\mathbb{N}}\subset{\mathcal{C}}^{1}$, such that $\lim_{n\to\infty}f_{n}=f$ in ${\mathcal{C}}^{\alpha-}.$ Therefore, by Lemma˜5.1, we have $\big(\int_{0}^{\cdot}\nabla f_{n}(X_{t})\mathrm{d}t\big)_{n\in\mathbb{N}}$ is a Cauchy sequence in $C_{p}^{1+(\alpha-1)H-}[0,1].$ We define ${\mathcal{H}}f=\lim_{n\rightarrow\infty}\int_{0}^{\cdot}\nabla f_{n}(X_{t})\mathrm{d}t$ in $C_{p}^{1+(\alpha-1)H-}[0,1].$
+Therefore, for $b\in{\mathcal{C}}^{\alpha}$, we can define $\int_{0}^{\cdot}\nabla b(X_{t})\mathrm{d}t:=\lim_{n\to\infty}\int_{0}^{\cdot}\nabla b_{n}(X_{t})\mathrm{d}t$ in probability.
+Moreover from Kolmogorov continuity criteria we know that it has a version which has $\theta$-Hölder continuous path for $\theta\in(0,1+(\alpha-1)H)\subset(0,1)$.
+Now we are at the position of showing Theorem˜2.4.
+Proof of Theorem˜2.4.
+Clearly Proposition˜5.2 shows the existence of $({\mathcal{H}}f)^{X}\in C_{p}^{1+(\alpha-1)H-}[0,1]$ which is bounded linear for $f\in{\mathcal{C}}^{\alpha}$, it also implies that the $c$ satisfying (2.7) for given $X$ is well-defined and $\|c\|_{{\mathcal{C}}^{1+(\alpha-1)H-}([0,1])}<\infty,{\mathbb{P}}$-a.s.
+Indeed from Kolmogorov continuity criteria and the property of Young integral we know that for any possible solution $c$, for any $k_{1}\in(\frac{1}{2},{1+(\alpha-1)H}),k_{2}\in(\frac{1}{2},(1+(\alpha-1)H)\wedge\alpha)=(\frac{1}{2},{1+(\alpha-1)H})$ so that $k_{1}+k_{2}>1$, for any $0\leqslant s\leqslant t\leqslant 1$ and $q>p$, we have
+$\|c(t)-c(s)\|_{L^{p}_{\omega}}\leqslant$ $\|\int_{s}^{t}c(r)\mathrm{d}({\mathcal{H}}b)_{r}^{X}\|_{L^{p}_{\omega}}+\frac{1}{2}\|b(X_{t})-b(X_{s})\|_{L^{p}_{\omega}}$
+$\leqslant C$ $\big\|\|c\|_{{\mathcal{C}}^{k_{1}}([0,1])}\|({\mathcal{H}}b)^{X}\|_{{\mathcal{C}}^{k_{2}}([0,1])}\big\|_{L^{p}_{\omega}}|t-s|^{k_{2}}+C\|b\|_{{\mathcal{C}}^{\alpha}({\mathbb{R}}^{d})}|t-s|^{\alpha}$
+$\leqslant C$ $\|c\|_{L_{\omega}^{q}({\mathcal{C}}^{k_{1}}([0,1]))}\|({\mathcal{H}}b)^{X}\|_{L_{\omega}^{\frac{qp}{q-p}}({\mathcal{C}}^{k_{2}}([0,1]))}|t-s|^{k_{2}}+C\|b\|_{{\mathcal{C}}^{\alpha}({\mathbb{R}}^{d})}|t-s|^{k_{2}}$
+$\leqslant C$ $\|b\|_{{\mathcal{C}}^{\alpha}({\mathbb{R}}^{d})}|t-s|^{k_{2}}.$ (5.2)
+Then, by Kolmogorov continuity criteria again we get $\mathbb{P}$}-a.s.
+$\|c\|_{{\mathcal{C}}^{1+(\alpha-1)H-}}<\infty$.
+Notice that (2.7) is a linear ODE in the sense of Young integral for given $X$, standard fixed point arguments gives us the existence and uniqueness of $c\in{\mathcal{C}}^{1+(\alpha-1)H-}([0,1])$ satisfying (2.7).
+Therefore we only need to show that (2.6) holds for such $c$ and for $b\in{\mathcal{C}}^{\alpha}$, $\alpha<1$.
+Denote $c^{n}:=n(X-X^{n})$.
+Theorem˜2.2 implies that $\sup_{n}\|c_{n}\|_{{\mathcal{C}}_{p}^{1+(\alpha-1)H-}[0,1]}<\infty,$ that is to say, for $\gamma<1+(\alpha-1)H$, from Kolmogorov continuity criteria, $\sup_{n}\|c_{n}\|_{L^{p}(\Omega,{\mathcal{C}}^{\gamma})}<\infty$, hence for ${\mathcal{L}}^{n}:=(X,X^{n},c^{n},c,c^{n}-c,B^{H})$, $\sup_{n}\|{\mathcal{L}}^{n}\|_{L^{p}(\Omega,({\mathcal{C}}^{\gamma})^{6})}<\infty.$
+By Sobolve embedding we know that there exists large $q$ and $\gamma_{1}<\gamma_{2}<\gamma$ so that compactly
+${\mathcal{C}}^{\gamma}([0,1],{\mathbb{R}}^{d})\subset\subset W_{q}^{\gamma_{2}}([0,1],{\mathbb{R}}^{d})\subset\subset{\mathcal{C}}^{\gamma_{1}}([0,1],{\mathbb{R}}^{d}),$ (5.3)
+hence the laws of $({\mathcal{L}}^{n})_{n\in{\mathbb{N}}}$ are tight on $W_{q}^{\gamma_{2}}([0,1],{\mathbb{R}}^{d})^{6}$ which is Polish and
+$\sup_{n}\||{\mathcal{L}}^{n}|\|_{L^{p}(\Omega,W_{q}^{\gamma_{2}}([0,1],{\mathbb{R}}^{d})^{6})}<\infty.$ (5.4)
+[…]
+[…]
+and
+$\hat{\mathbb{P}}-a.s.,\quad(\hat{\mathcal{L}}^{n})_{n\in{\mathbb{N}}_{1}}\quad\text{converges to }\quad{\mathcal{L}}\quad\text{ in }W_{q}^{\gamma_{2}}([0,1],{\mathbb{R}}^{d})^{6}.$ (5.6)
+Let $(\hat{\mathcal{F}}_{t})_{t\geqslant 0}$ be the augmentation of the filtration generated by $(\hat{X},\hat{X}^{n},\hat{B}^{H,n})$.
+(5.5), the continuity of $b$ and the strong well-posedness of (1.3) imply
+$\tilde{X}_{t}$ $=x_{0}+\int_{0}^{t}b(\tilde{X}_{s})\mathrm{d}s+\hat{B}_{t}^{H},\quad\tilde{X}_{t}^{k}=x_{0}+\int_{0}^{t}b(\tilde{X}_{s}^{k})\mathrm{d}s+\hat{B}_{t}^{H,k},$
+$\hat{X}_{t}^{k,n}$ $=x_{0}+\int_{0}^{t}b(\hat{X}_{k_{n}(s)}^{k,n})\mathrm{d}s+\hat{B}_{t}^{H,k},\quad\hat{c}^{n}=n(\tilde{X}^{n}-\hat{X}^{n,n}),$
+$\tilde{c}^{n}(t)$ $=\int_{0}^{t}\tilde{c}^{n}(s)\mathrm{d}({\mathcal{H}}b)^{\tilde{X}_{s}^{n}}+\frac{1}{2}(b(\tilde{X}_{t}^{n})-b(x_{0})),\quad t\in[0,1].$
+Following from (5.5) which implies the uniform integrability, (5.6) and Vitali’s theorem, for all $p\geqslant 1$ we have
+$\lim_{n\rightarrow\infty}\big\|\hat{\mathcal{L}}^{n}-{\mathcal{L}}\big\|_{L^{p}(\hat{\Omega},W_{q}^{\gamma_{2}}([0,1],{\mathbb{R}}^{d})^{6})}=0.$ (5.7)
+At this point we can also claim that $\tilde{X}=\hat{X} ^\hat{\mathbb{P}}$-a.s. from Theorem˜2.2 and the strong well-posedness of (1.3).
+Notice that for any $b\in{\mathcal{C}}^{\alpha}$, there exists a sequence $\{b_{n}\}_{n\in\mathbb{N}}\subset{\mathcal{C}}^{1}$ such that $\lim_{n\to\infty}b_{n}=b$ in ${\mathcal{C}}^{\alpha-}$, then for $m\in{\mathbb{N}}$
+$\hat{c}^{n}(t)=$ $n(\tilde{X}_{t}^{n}-\hat{X}_{t}^{n,n})=n\left(\int_{0}^{t}b(\tilde{X}_{s}^{n})\mathrm{d}s-\int_{0}^{t}b(\hat{X}_{k_{n}(s)}^{n,n})\mathrm{d}s\right)$
+$=$ $n\int_{0}^{t}(b-b_{m})(\tilde{X}_{s}^{n})-(b-b_{m})(\hat{X}_{k_{n}(s)}^{n,n})\mathrm{d}s+n\int_{0}^{t}b_{m}(\tilde{X}_{r}^{n})-b_{m}(\hat{X}_{r}^{n,n})\mathrm{d}s$
+$+n\int_{0}^{t}b_{m}(\hat{X}_{r}^{n,n})-b_{m}(\hat{X}_{k_{n}(s)}^{n,n})\mathrm{d}s=:J_{t}^{m,n}+K_{t}^{m,n}+L_{t}^{m,n}.$ (5.8)
+Following from Lemma˜4.2, Lemma˜4.3 and Lemma˜4.4 we get
+$\big\|\sup_{t\in[0,1]}|J_{t}^{m,n}|\big\|_{L_{\omega}^{p}}\lesssim\|b-b_{m}\|_{{\mathcal{C}}^{\alpha-}};$ (5.9)
+for $K^{m,n}$
+$K_{t}^{m,n}=\int_{0}^{t}\int_{0}^{1}\nabla b_{m}(\theta\tilde{X}_{r}^{n}+(1-\theta)\hat{X}_{r}^{n,n})\cdot\hat{c}_{r}^{n}\mathrm{d}\theta\mathrm{d}r,$
+since $^\hat{\mathbb{P}}$-a.s. $\hat{c}^{n}$ converges to $\hat{c}$, $\hat{X}^{n,n}$ converges to $\hat{X}$, and $\tilde{X}^{n}$ converges to $\tilde{X}=\hat{X}$, $^\hat{\mathbb{P}}$-a.s. we have
+$\lim_{n\rightarrow\infty}K_{t}^{m,n}=\int_{0}^{t}\nabla b_{m}(\hat{X}_{r})\cdot\hat{c}_{r}\mathrm{d}r=\int_{0}^{t}\hat{c}_{r}\mathrm{d}({\mathcal{H}}b_{m})_{r}^{\hat{X}},\quad\hat{\mathbb{P}}-a.s.$ (5.10)
+Concerning $L^{m,n}$, following from (1.9) and the fact a.s.
+$B^{H}\in{\mathcal{C}}^{H-\varepsilon}([0,1])$ for sufficiently small $\varepsilon>0$ by Kolmogorov continuity criteria, which means there exists a small enough $\varepsilon^{\prime}>0$ so that a.s.
+$B_{H}\in{\mathcal{C}}^{1+\varepsilon^{\prime}}$, we have a.s.
 (denote $\{c\}:=c-\lfloor c\rfloor$ for $c\in{\mathbb{R}}_{+}$)
-$n(X_{s}^{n}-X_{k_{n}(s)}^{n})$ $=n(s-k_{n}(s))b(X_{k_{n}(s)}^{n})+n(B_{s}^{H}-B_{k_{n}(s)}^{H})$
-$=\{ns\}b(X_{k_{n}(s)}^{n})+\{ns\}(B_{s}^{H})^{\prime}+o(n^{-\varepsilon^{\prime}});$
-together with the condition that $\nabla b$ is continuous and bounded, moreover $X_{r}^{n}\rightarrow X_{r}$ and $X_{k_{n}(s)}^{n}\rightarrow X_{s}$ a.s.
-from Theorem˜2.2, dominated convergence theorem shows as $n\rightarrow\infty$, we have
-$c^{n}(t)$ $-\int_{0}^{t}\exp\left(\int_{s}^{t}\nabla b(X_{r})dr\right)\nabla b(X_{s})\{ns\}(b(X_{s})+(B_{s}^{H})^{\prime})\mathrm{d}s\rightarrow 0\quad a.s.$ (5.2)
-Besides, observe
-$\int_{0}^{t}\exp\left(\int_{s}^{t}\nabla b(X_{r})dr\right)\nabla b(X_{s})\{ns\}(b(X_{s})+(B_{s}^{H})^{\prime})\mathrm{d}s$
-$\rightarrow\frac{1}{2}\int_{0}^{t}\exp\left(\int_{s}^{t}\nabla b(X_{r})dr\right)\nabla b(X_{s})(b(X_{s})+(B_{s}^{H})^{\prime})\mathrm{d}s\quad a.s.$
-since $\{n\cdot\}$ converges to $\frac{1}{2}$ weakly in $L^{2}([0,t])$ as $n\rightarrow\infty$, the above combining with (5.2) completes the proof. ∎
-Remark 5.2.
-Notice from Theorem˜5.1, we verify that indeed order $n$ reaches the optimal rate of EM (1.9), although for $b\in{\mathcal{C}}^{1}$ instead of assuming $b\in{\mathcal{C}}^{\alpha}$ for $\alpha<1$.
-Putting Theorem˜2.2 and Theorem˜5.1 together we can evidently say that order $n$ reaches the optimal (upper bound) rate of EM also for $b\in{\mathcal{C}}^{\alpha}$ simply by the embedding ${\mathcal{C}}^{1}\subset{\mathcal{C}}^{\alpha}$.
-However, concerning the lower bound of (1.9) for $\alpha<1$ we can not say too much so far.
+$n(\hat{X}_{r}^{n}-\hat{X}_{k_{n}(r)}^{n,n})$ $=n(r-k_{n}(r))b(\hat{X}_{k_{n}(r)}^{n,n})+n(\hat{B}_{r}^{H,n}-\hat{B}_{k_{n}(r)}^{H,n})$
+$=\{nr\}b(\hat{X}_{k_{n}(r)}^{n,n})+\{nr\}(\hat{B}_{r}^{H,n})^{\prime}+o(n^{-\varepsilon^{\prime}});$
+together with the condition that $\nabla b_{m}$ is continuous and bounded, moreover $\hat{X}_{r}^{n,n}\rightarrow\hat{X}_{r}$ and $\hat{X}_{k_{n}(r)}^{n,n}\rightarrow\hat{X}_{r}$ a.s.
+from Theorem˜2.2, dominated convergence theorem shows as $n\rightarrow\infty$, we have $^\hat{\mathbb{P}}$-a.s.
+$\lim_{n\rightarrow\infty}L_{t}^{m,n}$
+$=\lim_{n\rightarrow\infty}n\int_{0}^{t}\int_{0}^{1}\nabla b_{m}(\theta\hat{X}_{r}^{n,n}+(1-\theta)\hat{X}_{k_{n}(r)}^{n,n})\cdot(\hat{X}_{r}^{n,n}-\hat{X}_{k_{n}(r)}^{n,n})\mathrm{d}\theta\mathrm{d}r$
+$=\lim_{n\rightarrow\infty}\int_{0}^{t}\int_{0}^{1}\nabla b_{m}(\theta\hat{X}_{r}^{n,n}+(1-\theta)\hat{X}_{k_{n}(r)}^{n,n})\Big(\{nr\}\big(b(\hat{X}_{k_{n}(r)}^{n,n})+(B_{r}^{H})^{\prime}\big)+o(n^{-\varepsilon^{\prime}})\Big)\mathrm{d}r$
+$=\frac{1}{2}\int_{0}^{t}\nabla b_{m}(\hat{X}_{r})\big(b(\hat{X}_{r})+(\hat{B}_{r}^{H})^{\prime}\big)\mathrm{d}r=\frac{1}{2}\big(b_{m}(\hat{X}_{t})-b_{m}(x_{0})\big),$ (5.11)
+we have applied the fact that $\{n\cdot\}$ converges to $\frac{1}{2}$ weakly in $L^{2}([0,t])$ as $n\rightarrow\infty$ in deriving the penultimate equality and the lase equality holds due to elementary chain rule.
+Now we put (5.9), (5.10) and (5) together with taking $m,n\rightarrow\infty$ for both side of (5.8) then get $^\hat{\mathbb{P}}$-a.s.
+$\hat{c}(t)=$ $\int_{0}^{t}\hat{c}_{r}\mathrm{d}({\mathcal{H}}b)_{r}^{\hat{X}}+\frac{1}{2}(b(\hat{X}_{t})-b(x_{0})).$
+[…]
+Meanwhile following from the fact that $\hat{\mathbb{P}}$ a.s., $\tilde{X}^{n}$ converges to $\hat{X}$, $\tilde{c}^{n}$ converges to $\tilde{c}$ and the property of Young’s integral (2.1) we can conclude $\tilde{c}$ also satisfies (2.7) for such $\hat{X}$.
+It implies that $\hat{c}=\tilde{c}$ a.s.
+and $^\hat{\mathbb{P}}$-a.s., $\hat{c}^{n}-\tilde{c}^{n}$ converges to 0; hence we get that for given $X$, $c^{n}$ converges to $c$ in probability on $(\Omega,{\mathcal{F}},{\mathbb{P}})$.
+Appendix A Auxiliary Proofs
+Proof of (4) and (4.21).
+Similarly to (4) and (4.6), we have
+$\|A_{s,t}\|_{L^{p}_{\omega}}$ $=\|\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}f({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})$
+$\qquad-\mathcal{P}^{H}_{r-[s-(t-s)]}f({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi^{n}_{k_{n}(r)})\mathrm{d}r\|_{L^{p}_{\omega}}$
+$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\sup_{r\in[s,t]}\|\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}\|_{L_{\omega}^{p}}|t-s|^{1-(1-\alpha)H}$
+$\leqslant Cn^{-1}\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1-(1-\alpha)H},$ (A.1)
+where in the second inequality we used $\|\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}\|_{L_{\omega}^{p}}\leqslant\|b\|_{{\mathcal{C}}^{0}}n^{-1}$. Hence (4) holds.
+Next we verify (4.21). Let $(s,u,t)\in\overline{[0,1]}_{1}^{3}$. Recall the definition of $s_{i},i=1,\dots,6$ in (4.3). Similarly to (4.7), we can write
+${\mathbb{E}}^{s_{1}}\delta A_{s,u,t}=$ ${\mathbb{E}}^{s_{1}}\int_{s_{4}}^{s_{5}}\mathcal{P}_{r-s_{3}}^{H}f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-\mathcal{P}_{r-s_{3}}^{H}f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
+$\qquad\qquad-\mathcal{P}^{H}_{r-s_{3}}f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})+\mathcal{P}_{r-s_{3}}^{H}f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
+$+{\mathbb{E}}^{s_{1}}\int_{s_{5}}^{s_{6}}\mathcal{P}_{r-s_{2}}^{H}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-\mathcal{P}_{r-s_{2}}^{H}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
+$\qquad\qquad-\mathcal{P}_{r-s_{2}}^{H}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})+\mathcal{P}_{r-s_{3}}^{H}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
+$=$ $:I_{1}+I_{2}.$
+The two terms are treated in the exactly same manner, so we only detail $I_{1}$. Similarly to (4.8), we get
+$\|I_{1}\|_{L_{\omega}^{p}}\leqslant$ $C\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(1-\alpha)H}\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\big\|_{L_{\omega}^{p}}$
+$\qquad+(r-s_{3})^{-(2-\alpha)H}\big\||{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\cdot{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\big\|_{L_{\omega}^{p}}\mathrm{d}r.$ (A.2)
+Similar to (4.10), we have
+$\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\|_{L^{p}_{\omega}}\leqslant\|{\mathbb{E}}^{s_{1}}|(\varphi_{r}^{n}-\varphi_{r}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{r)}^{n})|\|_{L^{p}_{\omega}}.$ (A.3)
+We note that
+$\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}$ $=\int_{k_{n}(r)}^{r}b(B^{H}_{k_{n}(t)}+\varphi_{k_{n}(t)}^{n})\mathrm{d}t=(r-k_{n}(r))b(B_{k_{n}(r)}^{H}+\varphi^{n}_{k_{n}(r)})\in\mathcal{F}_{k_{n}(r)}.$
+When $s_{1}\in[k_{n}(r),r],$ we have
+$\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})=0;$
+[…]
+${\mathbb{E}}^{s_{1}}|(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})|$
+$=$ $(r-k_{n}(r)){\mathbb{E}}^{s_{1}}|b(B_{k_{n}(r)}^{H}+\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}b(B_{k_{n}(r)}^{H}+\varphi_{k_{n}(r)}^{n})|$
+$\leqslant$ $2(r-k_{n}(r)){\mathbb{E}}^{s_{1}}|b(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})-b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+\varphi_{s_{1}}^{n})|$
+$\leqslant$ $C(r-k_{n}(r)){\mathbb{E}}^{s_{1}}(|B_{k_{n}(r)}^{H}-{\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}|^{\alpha}+|\varphi_{k_{n}(r)}^{n}-\varphi_{s_{1}}^{n}|^{\alpha}).$
+Moreover, using (3.5) and (4.2), we have
+${\mathbb{E}}^{s_{1}}|(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})|$
+$\leqslant$ $C(r-k_{n}(r))(|k_{n}(r)-s_{1}|^{\alpha H}+|k_{n}(r)-s_{1}|^{\alpha})\leqslant\frac{C}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|r-s_{1}|^{\alpha}$
+where we used the fact $H>1$ in the second inequality. Plugging it into (A.3), we get
+$\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\|_{L^{p}_{\omega}}\leqslant\frac{C}{n}|r-s_{1}|^{\alpha}.$ (A.4)
+Meanwhile (4.9) and (4.2) yield
+${\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\leqslant C|r-s_{1}|^{1+\alpha H},$ (A.5)
+$\|{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})\|_{L_{\omega}^{p}}\leqslant C\|\varphi_{\cdot}^{n}-\varphi_{k_{n}(\cdot)}^{n}\|_{C_{p}^{0}}\leqslant\frac{C}{n}.$ (A.6)
+Applying (A.4), (A.5) and (A.6) into (A.2) gives us
+$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant C\frac{\|f\|_{{\mathcal{C}}^{\alpha}}}{n}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(1-\alpha)H}(r-s_{1})^{\alpha}\mathrm{d}r+(r-s_{3})^{-(2-\alpha)H}(r-s_{1})^{1+\alpha H}\mathrm{d}r$
+$\leqslant C\frac{\|f\|_{{\mathcal{C}}^{\alpha}}}{n}(t-s)^{(1+\alpha-(1-\alpha)H)\wedge(2+(2\alpha-2)H)}.$ (A.7)
+The same bound holds on $I_{2}$. Hence we get (4.21).
 Acknowledgment
@@
 12271030).
-We are grateful to Prof. Rongchan Zhu (Beijing Institute of Technology) for her valuable suggestions.
+We are grateful to Prof. Rongchan Zhu (Beijing Institute of Technology) for her fruitful advices. We also are grateful for valuable suggestion on the optimality result from an anonymous reviewer.